### v2

# SPIRAL: Self-Play on Zero-Sum Games Incentivizes Reasoning via Multi-Agent Multi-Turn Reinforcement Learning

Abstract
Recent advances in reinforcement learning have shown that language models can develop sophisticated reasoning through training on tasks with verifiable rewards, but these approaches depend on human-curated problem-answer pairs and domain-specific reward engineering. We introduce SPIRAL, a self-play framework where models learn by playing multi-turn, zero-sum games against continuously improving versions of themselves, eliminating the need for human supervision. Through self-play, SPIRAL generates an infinite curriculum of progressively challenging problems as models must constantly adapt to stronger opponents. To enable this self-play training at scale, We implement a fully online, multi-turn, multi-agent reinforcement learning system for LLMs and propose role-conditioned advantage estimation (RAE) to stabilize multi-agent training. Using SPIRAL, self-play on zero-sum games produces reasoning capabilities that transfer broadly. Training Qwen3-4B-Base on Kuhn Poker alone achieves 8.6% improvement on math and 8.4% on general reasoning, outperforming SFT on 25,000 expert game trajectories. Analysis reveals that this transfer occurs through three cognitive patterns: systematic decomposition, expected value calculation, and case-by-case analysis. Multi-game training (TicTacToe, Kuhn Poker, Simple Negotiation) further enhances performance as each game develops distinct reasoning strengths. Applying SPIRAL to a strong reasoning model (DeepSeek-R1-Distill-Qwen-7B) can still lead to 2.0% average improvement. These results demonstrate that zero-sum games naturally develop transferable reasoning capabilities, highlighting a promising direction for autonomous reasoning development.                          [Uncaptioned image] Spiral-RL/SPIRAL       [Uncaptioned image] Spiral-RL
Figure 1: SPIRAL achieves consistent improvements over base models across game performance and reasoning benchmarks. It also surpasses SFT on expert game trajectories and RL baselines trained against fixed opponents (Mistral and Gemini).
1 Introduction
Recent breakthroughs in language model reasoning, including OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek Team, 2024), reveal that reinforcement learning (RL) can unlock dramatic improvements in Chain-of-Thought reasoning (Wei et al., 2022). Through outcome-based rewards, RL enables models to develop generalizable reasoning strategies and consistently solve complex problems where supervised fine-tuning shows limited progress.
However, current approaches to reasoning enhancement face a fundamental scalability bottleneck: their dependence on carefully engineered reward functions, domain-specific datasets, and expert supervision (DeepSeek Team, 2024; Ouyang et al., 2022; Bai et al., 2022). Each new reasoning domain requires experts to craft evaluation metrics, curate training problems, and validate reasoning traces. This manual process becomes increasingly unsustainable as we pursue more general intelligence, limiting both the scale and diversity of reasoning challenges that models can learn from.
Self-play offers a solution to this scalability crisis by eliminating the need for human supervision in training data creation (Silver et al., 2017; Tesauro, 1995). In self-play, models learn by competing against copies of themselves, where each match outcome provides automatic feedback. As the model improves, its opponent improves equally, maintaining a consistent challenge level that drives continuous learning. This paradigm has already revolutionized AI across many domains: from TD-Gammon’s backgammon supremacy (Tesauro, 1995) to AlphaGo’s conquest of Go (Silver et al., 2016, 2017) to OpenAI Five’s mastery of complex team coordination (Berner et al., 2019). Yet despite these successes, applying self-play to enhance language model reasoning remains largely unexplored. Prior attempts have been limited to simple word games with offline updates (Cheng et al., 2024), LoRA adaptations that constrain learning capacity (Dettmers et al., 2023; Park et al., 2025), or single-turn code generation tasks (Zhao et al., 2025), falling short of leveraging multi-turn competitive dynamics that enable extended strategic reasoning.
We introduce SPIRAL, which applies self-play to two-player zero-sum language games for developing reasoning capabilities. Zero-sum language games provide ideal training environments: they require strategic thinking and planning while remaining simple enough to enable stable learning, with clear rules that make outcomes verifiable. SPIRAL offers two key advantages. First, unlike traditional RLVR approaches that depend on human-curated problem-answer pairs, it generates infinite training data through game dynamics alone. Second, compared to fixed-opponent training (see Fig. 2), self-play prevents overfitting to static strategies by continuously evolving the challenge level. However, implementing this vision for LLMs presents significant technical challenges. The computational demands of multi-turn, multi-agent autoregressive generation require sophisticated distributed systems, while standard RL algorithms suffer from high variance in multi-agent settings. To address these challenges, we implement a fully online, multi-turn, multi-agent reinforcement learning system with a distributed actor-learner architecture. We also introduce role-conditioned advantage estimation (RAE), which stabilizes training by normalizing rewards relative to each player’s expected performance. Without RAE, our ablation experiments show that models suffer from thinking collapse, progressively abandoning reasoning traces that are critical for generalization.
Key Findings. Training on zero-sum games produces reasoning capabilities that transfer broadly. SPIRAL trained solely on Kuhn Poker improves Qwen3-4B-Base by 8.6% on mathematical reasoning and 8.4% on general reasoning, outperforming SFT on 25,000 expert game trajectories despite using only game states with no mathematical content. This transfer works through three cognitive patterns we identify: systematic decomposition, expected value calculation, and case-by-case analysis. These patterns develop effectively through self-play, as fixed-opponent training fails while self-play continuously improves via adaptive challenge. Building on this, different games develop complementary skills (TicTacToe for spatial reasoning, Kuhn Poker for probabilistic thinking, and Simple Negotiation for strategic optimization) that combine synergistically in multi-game training. All these benefits require our Role-conditioned Advantage Estimation (RAE); without it, models abandon reasoning after 200 steps. The approach generalizes beyond weak models, with SPIRAL improving even strong models like DeepSeek-R1-Distill-Qwen-7B by 2.0%.
These results demonstrate that zero-sum games can serve as powerful reasoning gymnasiums, developing cognitive skills that transfer far beyond their training domain. Building on these findings, our work makes the following contributions:
Figure 2: From human-designed rewards to self-discovered reasoning through SPIRAL. Left: Traditional RL requires human experts to design complex reward functions. Middle: Fixed opponent training leads to exploitation of static strategies. Right: SPIRAL enables continuous reasoning improvement through self-play, where both players develop increasingly sophisticated strategies without human supervision.
  1. 1.
    A Fully Online, Multi-Turn, Multi-Agent RL Framework for LLMs: We develop a distributed actor-learner architecture that enables online self-play with full-parameter updates across multiple two-player zero-sum language games. The multi-turn aspect trains models to reason through sequential decisions, directly preparing them for complex multi-step problem solving. Unlike prior offline approaches, this provides a continuous curriculum as the model must adapt to an ever-improving opponent: itself. We release our implementation111 https://github.com/spiral-rl/spiral. to facilitate further research.
  2. 2.
    Role-conditioned Advantage Estimation (RAE): We introduce a variance-reduced advantage estimator specifically designed for multi-agent settings. By normalizing rewards relative to each player’s expected performance, RAE reduces variance in multi-agent learning. Our ablation studies empirically validate that RAE stabilizes learning and prevents thinking collapse. Without it, models progressively abandon reasoning traces, which is critical for generalization.
  3. 3.
    Empirical Discovery of Transfer: We demonstrate that self-play on zero-sum games improves both out-of-distribution game performance and academic reasoning benchmarks without domain-specific training data. Our analysis identifies reasoning patterns (systematic decomposition, expected value calculation, case-by-case analysis) that transfer from games to mathematics at measurable rates. Different games develop specialized skills that transfer to related domains, and multi-game training combines these abilities synergistically.
2 Related Work
Reinforcement Learning for LLM Reasoning. Reinforcement learning has evolved from alignment tasks using RLHF (Jaques et al., 2019; Ouyang et al., 2022; Bai et al., 2022) to directly improving reasoning capabilities. Recent models like OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek Team, 2024) demonstrate that RL with verifiable rewards (RLVR) can unlock chain-of-thought reasoning using rule-based rewards (Lightman et al., 2023; Uesato et al., 2022). However, these approaches depend on human-curated problem sets and domain-specific reward engineering. SPIRAL eliminates this dependency by using self-play games to generate infinite reasoning challenges without human supervision.
Multi-Agent RL for Language Models. Implementing MARL for full-scale LLMs presents significant technical challenges (Wan et al., 2025; Liu et al., 2025a). Prior work has made various compromises: Sarkar et al. (2025) uses RNNs instead of transformers; Jacob et al. (2022) focuses on simplified environments that don’t require full autoregressive generation; and Liao et al. (2024) shows the effectiveness of self-play with SFT on proprietary models. In contrast, SPIRAL implements fully online, full-parameter MARL through a distributed actor-learner architecture, enabling continuous adaptation to evolving opponents across multiple games.
Self-Play for Autonomous Improvement. Self-play in LLMs began with alignment methods like SPIN (Chen et al., 2024) and Self-Rewarding Language Models (Yuan et al., 2024). Recent work explores capability improvement: SPAG (Cheng et al., 2024) applies self-play to Adversarial Taboo but uses offline updates and remains confined to a single word game; SPC (Chen et al., 2025) and Genius (Xu et al., 2025) require human task distributions; Absolute Zero (Zhao et al., 2025) generates coding tasks but needs deterministic verification. SPIRAL demonstrates that self-play on strategic games improves academic benchmarks by 8.7% average without domain-specific data.
LLMs in Gaming. Games serve as testbeds for evaluating LLM capabilities (Paglieri et al., 2024; Ruoss et al., 2024; Zhang et al., 2024; Duan et al., 2024), but recent work explores games as training domains (Feng et al., 2024). LMRL-Gym (Abdulhai et al., 2023) benchmarks RL algorithms for LLMs in game environments; RAGEN (Wang et al., 2025) uses single-agent multi-turn RL; ViGaL (Xie et al., 2025) shows arcade games improve math reasoning; Divide-Fuse-Conquer (Zhang et al., 2025) trains on grouped games using offline learning. SPIRAL uniquely combines: (1) multi-agent self-play where both players share parameters, (2) fully online learning with continuous opponent evolution, and (3) transfer from zero-sum language games to reasoning without seeing any benchmark-related problems during training.
3 Preliminaries
Turn-level Markov Decision Process (MDP). Language model training traditionally formulates generation as a token-level MDP (Bellman, 1957; Rafailov et al., 2024) where each action is a single token from vocabulary $\mathcal{V}$. For multi-turn reasoning and game-playing, we instead adopt a turn-level MDP formulation $\mathcal{M}=(\mathcal{S},\mathcal{A},T,r,\gamma)$. Here, states $\mathcal{S}$ represent complete contexts (e.g., game configurations, problem states, or conversation histories), actions $\mathcal{A}$ are complete responses (containing many tokens), the transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ determines state dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides immediate rewards, and $\gamma\in[0,1]$ is the discount factor. The return is defined as the discounted sum of rewards: $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r_{t}$.
The key distinction: in token-level MDPs, each decision outputs one token; in turn-level MDPs, each decision produces a complete multi-token response before transitioning. At each turn $t$, the language model observes state $s_{t}$ and generates:
$y_{t}=\langle\text{think}\rangle c_{t}\langle/\text{think}\rangle\langle\text{answer}\rangle a_{t}\langle/\text{answer}\rangle,$ (1)
where $c_{t}$ externalizes reasoning and $a_{t}\in\mathcal{A}$ is the executable action. (See App. A for how existing SFT and RLVR paradigms adapt to turn-level MDPs.)
Two-Player Zero-Sum Markov Games. We extend the single-agent MDP to competitive settings with a two-player zero-sum Markov game (Littman, 1994) $\mathcal{G}=(\mathcal{S},\mathcal{A}_{0},\mathcal{A}_{1},T,r,\gamma)$, where $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are the action spaces for player 0 and player 1 respectively. The zero-sum property requires:
$r_{0}(s,a^{(0)},a^{(1)})+r_{1}(s,a^{(0)},a^{(1)})=0\quad\forall s,a^{(0)},a^{(1)},$ (2)
where $a^{(0)}\in\mathcal{A}_{0}$ and $a^{(1)}\in\mathcal{A}_{1}$ denote actions taken by each player. Given trajectory $\tau=\{(s_{t},a_{t}^{(0)},a_{t}^{(1)})\}_{t=0}^{T}$, the returns satisfy $R_{1}(\tau)=-R_{0}(\tau)$.
4 The SPIRAL Framework
Figure 3: The SPIRAL Framework. SPIRAL employs an actor-learner architecture for scalable self-play training. Parallel actors sample trajectories from a diverse set of games using vectorized environments. A single policy $\pi_{i}$ plays both roles, generating zero-sum, sparse reward game trajectories. The centralized learner processes these trajectories using Role-conditioned Advantage Estimation (RAE) to compute separate advantages, $A_{0}(s,a)$ and $A_{1}(s,a)$, for each role. These are then used for on-policy reinforcement learning updates.
We present SPIRAL, a framework that enables language models to develop generalizable reasoning capabilities through multi-turn competitive self-play. The framework is illustrated in Fig. 3.
4.1 SPIRAL as Self-Play on Turn-Based Zero-Sum Games
SPIRAL implements self-play through turn-based zero-sum language games from a collection $\mathcal{G}=\{G_{1},G_{2},...,G_{n}\}$. Each game $G_{i}$ is a two-player zero-sum Markov game, with the specific structure that players alternate turns rather than acting simultaneously. SPIRAL leverages three properties of these games:
First, the multi-turn structure mirrors sequential reasoning problems. Players alternate turns: at time $t$, player $p=t\bmod 2$ acts while the opponent waits, formally: $a_{t}^{(p)}\in\mathcal{A}_{p}$ and $a_{t}^{(1-p)}=\emptyset$. Unlike single-turn bandit settings, this framework trains models to maintain context, plan ahead, and adapt strategies based on evolving game states.
Second, the zero-sum competitive dynamics create pressure for continuous improvement. Each game uses sparse terminal rewards: $r_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})=0$ for all non-terminal states. At terminal state $s_{T}$, players receive opposite rewards:
$R_{0}(\tau)=\rho_{i}(s_{T}),\quad R_{1}(\tau)=-\rho_{i}(s_{T}),$ (3)
where $\rho_{i}:\mathcal{S}_{i}^{\text{terminal}}\rightarrow\{-1,0,1\}$ determines the outcome. This opposition forces models to develop robust strategies that cannot be easily exploited.
Finally, both players share a single policy $\pi_{\theta}$ with role conditioning. To enable this in the LLM setting, we prepend role-specific system prompts that specify whether the model is playing as Player 0 or Player 1 (see Sec. B.1 for example observations). At each turn, the active player generates:
$y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i}),$ (4)
where conditioning on $p$ is achieved through the system prompt [^2] We note that both $p$ and $G_{i}$ can be subsumed into state $s_{t}$, but we choose to make them explicit here for clarity.. From $y_{t}^{(p)}$, we extract the action $a_{t}^{(p)}$ to update the game state. This shared-parameter approach ensures that as the model improves at one role, it simultaneously faces a stronger opponent, creating an automatic curriculum.
4.2 Multi-Turn Self-Play Training Objective
In self-play, we could train two separate policies $\pi_{\theta_{0}}$ and $\pi_{\theta_{1}}$ for each player. Each would maximize their own expected return:
$J_{0}(\theta_{0})$ $=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta_{0}}\times\pi_{\theta_{1}}|G}\left[R_{0}(\tau)\right]$ (5)
$J_{1}(\theta_{1})$ $=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta_{0}}\times\pi_{\theta_{1}}|G}\left[R_{1}(\tau)\right]$ (6)
However, SPIRAL uses a single shared policy $\pi_{\theta}$ for both players, setting $\theta_{0}=\theta_{1}=\theta$. While this creates opposing objectives since $R_{1}(\tau)=-R_{0}(\tau)$ in zero-sum games, the role conditioning through system prompts enables the model to learn distinct strategies for each position within a unified policy framework.
To optimize this shared policy, we apply policy gradient methods. Using REINFORCE (Williams, 1992), the gradient becomes:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta}\times\pi_{\theta}|G}\left[\underbrace{\sum_{t\in T_{0}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(0)}|s_{t},0,G)\cdot R_{0}(\tau)}_{\text{Player 0 gradient}}+\underbrace{\sum_{t\in T_{1}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(1)}|s_{t},1,G)\cdot R_{1}(\tau)}_{\text{Player 1 gradient}}\right],$ (7)
where $T_{p}=\{t:t\bmod 2=p\}$ denotes turns where player $p$ acted. This formulation uses Monte Carlo returns which suffer from high variance, particularly problematic in self-play where the opponent’s strategy continuously evolves, making the environment highly non-stationary.
Algorithm 1 SPIRAL: Role-Balanced Multi-Turn Self-Play
1:Policy $\pi_{\theta}$, Games $\mathcal{G}=\{G_{1},...,G_{n}\}$, decay rate $\alpha\in[0,1]$
2:Initialize baselines $b_{G_{i},p}=0$ for all $G_{i}\in\mathcal{G}$, $p\in\{0,1\}$
3:while not converged do
4:     // Self-Play Trajectory Collection
5:     $\mathcal{B}\leftarrow\emptyset$
6:     for $k=1$ to $K$ actors in parallel do
7:         Sample game $G_{i}\sim\mathcal{G}$, initialize $s_{0}\sim G_{i}$
8:         for turn $t=0,1,2,...$ until terminal do
9:              $p\leftarrow t\bmod 2$triangleright Determine active player
10:              $y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i})$triangleright Generate reasoning + action
11:              $a_{t}^{(p)}\leftarrow\text{extract\_action}(y_{t}^{(p)})$
12:              $a_{t}^{(1-p)}\leftarrow\emptyset$triangleright Inactive player
13:              $s_{t+1}\leftarrow T_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})$
14:         end for
15:         $R_{0}\leftarrow\rho_{i}(s_{T})$, $R_{1}\leftarrow-R_{0}$
16:         Add $(\tau,G_{i})$ to batch $\mathcal{B}$triangleright Store trajectory with its game
17:     end for
18:     // Role-Balanced Policy Optimization
19:     for $(\tau,G_{i})\in\mathcal{B}$ do
20:         for $p\in\{0,1\}$ do
21:              $b_{G_{i},p}\leftarrow\alpha b_{G_{i},p}+(1-\alpha)R_{p}(\tau)$
22:              $A_{G_{i},p}(\tau)\leftarrow R_{p}(\tau)-b_{G_{i},p}$
23:         end for
24:     end for
25:     Update $\theta$ using policy gradient with advantages $A_{G_{i},p}$ (Eq. 10)
26:end while
4.3 Policy Optimization with Role-conditioned Advantage Estimation
To reduce the high variance inherent in multi-agent REINFORCE, we introduce Role-conditioned Advantage Estimation (RAE). In two-player games, even with a shared policy, different roles may have different expected returns due to game asymmetries (e.g., first-move advantage in TicTacToe, information asymmetry in Kuhn Poker).
RAE maintains separate baselines $b_{G,p}$ for each game $G\in\mathcal{G}$ and role $p\in\{0,1\}$, where each baseline estimates the expected return $\mathbb{E}[R_{p}(\tau)]$ for that role in that game. We update these baselines using exponential moving average (EMA) with a decay rate $\alpha\in[0,1]$:
$b_{G,p}\leftarrow\alpha b_{G,p}+(1-\alpha)R_{p}(\tau)$ (update baseline) (8)
$A_{G,p}(\tau)=R_{p}(\tau)-b_{G,p}$ (compute advantage) (9)
This provides better variance reduction than a global baseline by accounting for role-specific asymmetries, such as the first-move advantage. The variance-reduced policy gradient becomes:
$\nabla_{\theta}J_{\text{SPIRAL}}(\theta)=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta}\times\pi_{\theta}|G}\left[\sum_{p\in\{0,1\}}\sum_{t\in T_{p}}A_{G,p}(\tau)\cdot\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(p)}|s_{t},p,G)\right]$ (10)
By centering returns around role-specific expectations, RAE ensures that gradient updates reflect genuine learning signal rather than inherent advantages of certain player positions. Note that we do not normalize by response length to avoid length bias (Liu et al., 2025b). The complete procedure is in Algorithm 1.
4.4 Implementation
To implement SPIRAL, we develop a truly online multi-agent, multi-turn RL system for finetuning LLMs. Our training framework builds on Oat (Liu et al., 2024), which provides interfaces of a distributed actor-learner architecture (Espeholt et al., 2018). We instantiate actors to execute the self-play loop, using vLLM (Kwon et al., 2023) for efficient model inference and TextArena (Guertler et al., 2025) to simulate the language games. The resulting multi-turn, multi-game self-play experiences are used to update the LLM via policy gradient methods (Sutton and Barto, 2018), incorporating our proposed Role-conditioned Advantage Estimation in the collocated learner.
5 Experiment Setup
We design experiments to carefully test how training on different zero-sum games can improve reasoning capabilities that generalize to novel downstream math and reasoning tasks, as well as novel zero-sum games. This section introduces the selection of game environments (Sec. 5.1) and the training (Sec. 5.2) and evaluation (Sec. 5.3) settings.
We investigate four key questions:
  1. 1.
    RQ1: Can self-play on zero-sum games improve math and general reasoning capabilities? We test whether competitive environment alone can develop transferable reasoning capabilities without domain-specific training data, and analyze emergent patterns to understand transfer mechanisms.
  2. 2.
    RQ2: Does self-play’s automatic curriculum outperform fixed-opponent training? We evaluate whether the automatic curriculum with adaptive difficulty generated through self-play produces more robust reasoning than training against static opponents.
  3. 3.
    RQ3: Do different games develop specialized reasoning skills? We test whether distinct game mechanics cultivate complementary cognitive abilities and whether multi-game training creates synergistic benefits.
  4. 4.
    RQ4: Is Role-Conditioned Advantage Estimation (RAE) essential for stable self-play training? We ablate our variance reduction technique to determine whether RAE is necessary to stabilize training and prevent thinking collapse in zero-sum self-play with shared parameters.
5.1 Game Environments
We select three games from TextArena (Guertler et al., 2025) with distinct cognitive requirements:
TicTacToe (Spatial Reasoning) (Beck, 2008). A 3×3 grid game requiring pattern recognition and adversarial planning. Players must identify winning configurations, block opponent threats, and plan multi-step forcing sequences. We hypothesize these skills transfer to geometric problem-solving and spatial visualization tasks. Tic-Tac-Toe is a deterministic, perfect-information game, which isolates pure strategic reasoning from uncertainty management.
Kuhn Poker (Probabilistic Reasoning) (Kuhn, 1950). A minimal poker variant with three cards (Jack, Queen, King) where players bet with hidden information. Success requires probability calculation, opponent modeling, and decision-making under uncertainty. We expect these capabilities to transfer to problems involving probability, expected value, and strategic uncertainty.
Simple Negotiation (Strategic Optimization) (Nash, 1950). A resource trading game where two players exchange Wood and Gold with opposing valuations to maximize portfolio value. Success requires multi-step planning, theory of mind to model opponent preferences, and strategic communication through proposals and counteroffers. We hypothesize these skills transfer to optimization problems, resource allocation tasks, and multi-constraint reasoning that requires balancing competing objectives. Please see Fig. 7 in App. B for example environment observations.
5.2 Training Settings
We employ Qwen3-4B-Base (Yang et al., 2025) as the base model for all experiments. The training spans 400 steps with 128 samples per step, totaling 51,200 game transitions. We use a sampling temperature $\tau=1.0$ to collect game experiences and Adam (Kingma and Ba, 2014) with a constant learning rate of $1\times 10^{-6}$ to optimize the model. The training batch size is set as 128 and the discount factor $\gamma=1.0$ (no discounting). We use $\alpha=0.95$ for the exponential moving average decay of role-specific baselines. Detailed hyperparameters are listed in Table 7 in App. B. Each experiment takes around 25 hours for full parameter tuning on 8 H100 GPUs with our distributed actor-learner architecture.
5.3 Evaluation Metrics
We track three key metrics throughout the training process: the frequency of invalid moves, the game lengths and the win rates against Gemini-2.0-Flash-Lite [^3] Accessed via https://openrouter.ai/google/gemini-2.0-flash-lite-001., to monitor the learning dynamics and determine the best models. To evaluate the trained models, we test their win rates against a fixed opponent on both training and unseen games and standard reasoning benchmarks, as detailed below.
Out-of-Distribution Generalization to Unseen Games. We evaluate transfer learning on seven unseen games that test whether learned skills generalize beyond training domains: Snake and Connect Four (spatial reasoning), Pig Dice and Liar’s Dice (probabilistic reasoning), Truth and Deception (strategic optimization). These games specifically probe whether spatial reasoning from TicTacToe, probabilistic reasoning from Kuhn Poker, and strategic optimization from Simple Negotiation transfer to novel game mechanics. We report the average win rate of our models against Gemini-2.0-Flash [^4] Accessed via https://openrouter.ai/google/gemini-2.0-flash-001. across 512 independent games, with the starting player randomized in each match to eliminate any first-move advantage.
Transfer to Standard Reasoning Benchmarks. To investigate whether the reasoning abilities developed through gameplay could transfer to non-game contexts, we evaluate our models on a suite of established benchmarks. For mathematical reasoning, we use MATH500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME24, AIME25, and AMC23 datasets, which cover a wide range of topics including algebra, geometry, and competitive mathematics. For general reasoning, we utilize GPQA (Rein et al., 2024), which consists of graduate-level science questions, and MMLU-Pro (Wang et al., 2024), a benchmark for multidisciplinary knowledge. All evaluations on these benchmarks are conducted in a zero-shot setting [^5] Except for the base model, for which we provide few-shot examples that follow the Qwen3 Report settings. to determine if game-induced reasoning could be successfully transferred to general problem-solving. We use sampling temperature $0.6$, top-p $0.95$ and $4$ rollouts for all evaluations.
Reasoning Pattern Analysis. We analyze the cognitive patterns that emerge during training and their transfer to mathematical reasoning. Using GPT-4.1 (gpt-4.1-2025-04-14(OpenAI, 2025) as LLM-as-Judge (Zheng et al., 2023), we classify reasoning traces from 290 game trajectories and 46,792 math problem solutions into three core patterns: Case-by-Case Analysis (systematic enumeration of scenarios), Expected Value Calculation (probabilistic decision-making), and Pattern Recognition (identifying regularities and structures). The complete prompts used for pattern discovery, classification, and transfer analysis are provided in App. D. We track the evolution of these patterns across training checkpoints (early: step 0, mid: step 128, late: step 400) in both game and math domains to measure transfer rates. This analysis reveals which game-induced reasoning strategies generalize to academic problem-solving and explains the mechanism behind cross-domain transfer.
6 Experimental Results and Findings
6.1 RQ1: Can self-play on zero-sum games improve math and general reasoning capabilities?
Table 1 presents our main results. SPIRAL, trained exclusively on Kuhn Poker through self-play, achieves notable improvements on both mathematical and general reasoning benchmarks. Notably, it even surpasses a model that was fine-tuned on 25,000 winning game trajectories from Qwen3-32B playing against itself.
Table 1: Reasoning benchmark performance. SPIRAL improves reasoning without any domain-specific training data.
Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average
Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 32.9
SFT (Qwen3-32B Distill) 76.0 16.7 13.3 38.4 50.0 31.2 33.0 48.8 39.7
SPIRAL (Ours) 76.4 13.3 10.0 38.4 57.5 42.4 37.0 57.7 41.6
Improvement +10.6 +3.3 +6.7 +5.1 +7.5 +18.1 +6.4 +10.5 +8.7
Few-shot evaluation following Qwen3 technical report.
The results show consistent improvements across mathematical reasoning (+10.6% MATH500, +6.7% AIME’25, +7.5% AMC-23), and general reasoning (+6.4% GPQA, +10.5% MMLU-Pro). Notably, SPIRAL achieves these gains despite never seeing a single mathematical equation, academic problem, or domain-specific training example during its entire training process. The 18.1% improvement on Minerva Math and overall 8.7% gain across math benchmarks demonstrate that competitive game environments successfully force discovery of reasoning patterns that transfer broadly across domains.
6.1.1 Case Study: How Game Strategies Transfer to Mathematical Reasoning
To understand this surprising transfer, we use GPT-4.1 (gpt-4.1-2025-04-14) to analyze 290 game trajectories and 46,792 math solutions across training checkpoints, identifying three core reasoning patterns that emerge during gameplay and transfer to math reasoning domains.
Pattern Discovery and Examples. We first illustrate what these patterns look like in both game and mathematical contexts:
Table 2: Representative reasoning patterns and their manifestation in games and mathematics. We use an LLM-as-a-judge framework to automatically evaluate the frequency of these patterns over training.
Pattern Game Context (Kuhn Poker) Math Transfer
Case-by-Case Analysis <think>
Case 1 - Fold: You lose 1 chip. You have 2 chips now.
Case 2 - Call: You have 0% chance of winning, so you will have 1 chip if you lose.
Since losing 1 chip (fold) is better than losing 2 chips (call and lose), the best action is to fold.
</think>
<think> To solve $m^{2}+2\cdot 3^{n}=m(2^{n+1}-1)$:
Case $n=0$: $m^{2}+2=m$, no real solutions
Case $n=1$: $m^{2}+6=3m$, no real solutions
Case $n=2$: $m^{2}+18=7m$, no real solutions
Case $n=3$: $m^{2}+54=15m$ gives $m=6,9$ (valid)
Case $n\geqslant 4$: No solutions (exponential dominance) </think>
Expected Value Calculation <think> With card J against opponent’s bet:
EV(call) = P(win) $\times$ 2 - P(lose) $\times$ 2
= 0 $\times$ 2 - 1 $\times$ 2 = -2
EV(fold) = -1 (lose ante)
Since EV(fold) > EV(call), fold is optimal. </think>
<think> For average of $A+BC+DEF$ over permutations:
E[A] = $\frac{\sum a_{i}}{6}=\frac{-2}{6}=-\frac{1}{3}$
E[BC] = $(E[B])(E[C])=(-\frac{1}{3})^{2}=\frac{1}{9}$
E[DEF] = $(E[D])(E[E])(E[F])=(-\frac{1}{3})^{3}=-\frac{1}{27}$
Total = $-\frac{1}{3}+\frac{1}{9}-\frac{1}{27}=-\frac{7}{27}$ </think>
Pattern Recognition <think> Previous rounds: Player 0 had K both times and won. Pattern identified: Player 0 likely has strong cards or bluffs consistently. Given I have J (weak), betting might exploit their aggressive calling pattern. </think> <think> Sum of three consecutive integers $(n-1)+n+(n+1)=3n$.
Pattern recognized: sum is always divisible by 3.
For perfect cube: $3n=k^{3}$, so $k$ must be divisible by 3.
Smallest: $k=3\Rightarrow 3n=27$ </think>
Pattern Evolution and Transfer Analysis. Using GPT-4.1 to classify reasoning strategies, we tracked how these patterns develop during training and transfer to mathematics:
Figure 4: Evolution of reasoning patterns during SPIRAL training and their transfer to mathematical reasoning. We track three core reasoning patterns (Pattern Recognition, Expected Value Calculation, and Case-by-Case Analysis) across 290 game trajectories and 46,792 math solutions. Left: In the game domain, all patterns show substantial growth, with Expected Value Calculation reaching 78% by late training. Middle: These patterns transfer to mathematical reasoning with varying effectiveness: Case-by-Case Analysis maintains high transfer (72% to 71%), Pattern Recognition shows amplification (35% to 45%), while Expected Value Calculation transfers more selectively (78% to 28%). Right: Math benchmark scores improve from 31.2 to 39.6 as these reasoning patterns develop, demonstrating that game-learned strategies enhance mathematical problem-solving capabilities.
Fig. 4 shows how reasoning patterns evolve in both game and math domains across training checkpoints. These patterns develop at different rates and transfer with varying effectiveness:
Case-by-Case Analysis: We find that the meta-cognitive skill of systematic enumeration, or breaking problems down into cases, transfers highly effectively from strategic games to math problems (as shown in Fig. 4). We hypothesize this is because it represents a domain-agnostic way of structuring thinking that generally improves reasoning performance. Whether analyzing opponent possibilities in Poker or solution branches in mathematics, the core skill is the same.
Expected Value Calculation: We find that training on games increases the use of expected value calculation from 15% to 78% in the games themselves, and from 12% to 28% in the math domain. While game-specific probabilistic reasoning is more frequent in games than in the math benchmark (likely because most math problems lack explicit decision-theoretic structure), we find that this form of reasoning benefits probability and optimization problems where EV concepts directly apply.
Pattern Recognition: Interestingly, this form of reasoning shows an amplification effect, where the proportion evident in the math domain actually exceeds that of the game domain. We hypothesize that because mathematics inherently requires pattern recognition, game training enhances an already-present mathematical skill which is then deployed even more frequently when the model is prompted to answer a math problem. The combined training produces stronger pattern recognition (45%) than games alone require (35%).
Transfer Mechanism. We believe three factors enable this cross-domain transfer:
1. Competitive pressure strips away memorization: Self-play opponents continuously evolve, forcing models to develop genuine reasoning rather than pattern matching.
2. Games isolate pure reasoning: Without domain complexity, games teach fundamental cognitive operations (enumeration, evaluation, synthesis) that generalize effectively.
3. Structured output bridges domains: The <think> format learned in games provides a reasoning scaffold that models reuse for mathematical problems.
Finding 1: Zero-sum games teach transferable reasoning without domain data Self-play on Kuhn Poker improves math and general reasoning benchmarks despite never seeing benchmark related problems. This works because gameplay forces discovery of three reasoning patterns: Case-by-Case Analysis, Expected Value Calculation, and Pattern Recognition. These skills are developed during competing against an evolving opponent.
6.2 RQ2: Does self-play’s automatic curriculum outperform fixed-opponent training?
Although self-play promises potentially infinite curriculum generation, it remains unclear whether the evolving challenge actually produces better learning than simply training against fixed opponents. We test whether adaptive difficulty through self-play leads to more robust reasoning development.
Experimental Setup. We compare four training settings on the Kuhn Poker environment: (1) Self-Play: the model trains against continuously updating copies of itself; (2) Random Opponent: the opponent always provides valid actions but poses minimal strategic challenge; (3) Mistral Opponent: using Mistral-Small-3 (Mistral, 2025) as a fixed intermediate-level opponent; and (4) Gemini Opponent,: using Gemini-2.0-Flash-Lite (Gemini Team, Google, 2025) as a fixed strong opponent. All variants use identical hyperparameters, infrastructure, and 400 training steps.
Results Analysis.
Figure 5 reveals two distinct failure modes in fixed-opponent training:
Figure 5: Performance comparison of self-play training and fixed-opponent baselines. All evaluations are averaged over multiple games/benchmarks (see Sec. 5.3). Mistral Opponent refers to against Mistral-Small-3; Gemini Opponent refers to against Gemini-2.0-Flash-Lite.
1. Curse of Turns in Format Learning: Training against random opponents leads to complete collapse. Although these rule-based random agents offer no strategic challenge, they always produce valid moves. Consequently, the model must generate correctly formatted, valid actions at every turn throughout the trajectory in order to receive any positive reward. However, the probability of producing a fully valid trajectory decreases exponentially with the episode length, making it extremely difficult to explore and learn the correct action format over long horizons.
2. Exploitation of Static Strategies: Fixed model-based opponents (Mistral, Gemini) facilitate format learning more effectively than random opponents, as they occasionally produce invalid moves that allow the agent to win and receive reward signals. However, we find that training against these static opponents often leads to overfitting to their fixed strategies, resulting in poor generalization to math and general reasoning benchmarks (middle and right of Fig. 5).
Continuous Adaptation Through Self-Play. Self-play uniquely maintains learning momentum throughout training and generalizes better because self-play automatically adjusts difficulty. As shown in Table 3, training against a fixed opponent (Gemini) initially gets zero win rate (thus no positive reward), leading to no effective learning before step 128 (see left of Fig. 5); at step 384 the agent learns to outperform the fixed opponent with $62.5\%$ win rate, gradually exploiting its static strategies. On the other hand, against previous versions of itself, self-play maintains 50-52% win rates throughout training, confirming the opponent evolves to match current capabilities.
Table 3: Win rates at different training stages of Gemini Opponent and Self-Play vs its opponent.
Training Stage Gemini Opponent Win Rate vs Gemini-2.0-Flash-Lite Self-Play Win Rate vs Self (t-16)
Step 16 0.0% 52.3%
Step 128 37.5% 51.7%
Step 384 62.5% 50.9%
The reasoning transfer results are even more striking. Self-play achieves 40% on math reasoning and 45% on general reasoning, outperforming the best fixed opponent (Gemini) by 5 and 3 percentage points respectively. The relative improvement demonstrates that the diverse strategies required by an evolving opponent create more generalizable reasoning patterns than exploiting static weaknesses.
Finding 2: Self-play creates adaptive curriculum that outperforms fixed opponents Self-play achieves better reasoning transfer than any fixed-opponent training. Rule-based random opponents cause complete collapse due to the curse of turns in format learning. Training against model-based opponents plateaus once the agent finds winning counter-strategies. Self-play avoids both failures by continuously adjusting difficulty: the opponent improves as the agent improves, forcing ongoing adaptation rather than exploitation of static weaknesses.
6.3 RQ3: Do different games develop specialized reasoning skills?
We investigate whether distinct game mechanics cultivate different cognitive abilities. If games truly serve as reasoning gymnasiums, then training on TicTacToe should develop different skills than training on Kuhn Poker or Simple Negotiation. Furthermore, we test whether combining multiple games produces synergistic benefits beyond the sum of individual training.
Experimental Design.
To isolate game-specific skill development, we train separate models exclusively on each game using identical hyperparameters and training duration. We then evaluate these specialists in three ways: (1) performance on their training game, (2) transfer to out-of-distribution games, and (3) generalization to math and general reasoning benchmarks. This design reveals whether games develop narrow tactics or transferable cognitive capabilities. For game performance evaluation, we use two complementary approaches: head-to-head competition between specialists to directly compare learned strategies, and performance against a fixed strong opponent (Gemini-2.0-Flash-Lite) to measure absolute skill levels. This dual evaluation reveals both relative strengths between specialists and their absolute capabilities.
Head-to-Head Competition Reveals Specialized Strategies.
We first examine how specialists perform when competing directly against each other. This head-to-head evaluation uses games that isolate specific cognitive skills: we pair each training game with an unseen game requiring similar abilities (Snake for spatial reasoning like TicTacToe, Pig Dice for probabilistic reasoning like Kuhn Poker, and Truth and Deception for strategic optimization like Simple Negotiation).
Table 4: Game specialists excel at both their training games and unseen games requiring similar cognitive skills. Each cell shows the win rate in head-to-head competition between specialists (e.g., 57.5% means TicTacToe specialist wins 57.5% of games against the other two specialists on TicTacToe). Bold indicates best performance in each column.
Training Games OOD Games (Similar Skills)
Model TicTacToe Kuhn Poker Simple Negotiation Snake Pig Dice Truth and Deception
(Spatial) (Probabilistic) (Strategic)
TicTacToe Specialist 57.5% 45.1% 30.4% 56.0% 56.7% 48.7%
Poker Specialist 45.5% 64.2% 37.7% 42.5% 91.7% 45.4%
Negotiation Specialist 40.5% 40.2% 62.7% 41.0% 1.1% 55.8%
Models trained exclusively on TicTacToe excel at spatial pattern recognition, achieving 57.5% win rate against other specialists on their training game. When tested on Snake, an unseen spatial reasoning game, they maintain 56.0% performance, demonstrating robust transfer. Kuhn Poker specialists develop probabilistic reasoning, dominating their training game (64.2%) and remarkably achieving 91.7% on Pig Dice. Simple Negotiation training produces models skilled in strategic optimization (62.7% on training game, 55.8% on Truth and Deception).
Multi-Game Training Creates Synergistic Benefits.
While head-to-head competition reveals specialized strategies, we also evaluate absolute performance against a strong fixed opponent (Gemini-2.0-Flash-Lite) across both training games and randomly sampled out-of-distribution games. The multi-game model demonstrates broader capabilities that often exceed specialist performance:
Table 5: Multi-game training achieves competitive performance across all training games while excelling at novel composite challenges. All win rates shown are against Gemini-2.0-Flash-Lite as a fixed opponent. The multi-game model outperforms all specialists on average, demonstrating that diverse game training develops more flexible reasoning.
Training Games OOD Games
Model TicTacToe KuhnPoker Simple Negotiation Liar’s Dice Connect Four Snake Average
Base Model 17.5% 21.5% 15.6% 16.7% 20.5% 20.6% 18.7%
Single-Game Specialists
      TicTacToe Specialist 56.6% 24.4% 30.5% 31.4% 30.8% 30.7% 34.1%
      Kuhn Poker Specialist 31.0% 48.5% 28.7% 24.9% 39.1% 33.9% 34.4%
      Simple Negotiation Specialist 27.7% 16.8% 39.1% 12.3% 32.0% 34.8% 27.1%
Multi-Game Model 54.3% 53.9% 33.2% 51.4% 39.3% 37.2% 44.9%
The power of multi-game training becomes evident in performance on novel games. On Liar’s Dice, individual specialists struggle, achieving only 24.9% (Kuhn Poker) and 12.3% (Simple Negotiation) against Gemini-2.0-Flash-Lite. However, the multi-game model achieves 51.4%, demonstrating emergent capabilities beyond any single training game. This pattern of multi-game superiority holds across the out-of-distribution games, with the multi-game model achieving the highest average performance (44.9%) compared to all specialists.
Multi-Game Training Enhances Even Strong Reasoning Models.
To test whether game-based skill development is limited to weaker models, we apply multi-game SPIRAL training to DeepSeek-R1-Distill-Qwen-7B, a model that already achieves 59.7% average on reasoning benchmarks (see Sec. C.2 for comprehensive results across all training configurations):
Table 6: Multi-game SPIRAL training improves reasoning capabilities across model scales. Even strong reasoning models benefit from game-based skill development.
Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average
Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 33.1
+ SPIRAL (Multi-Game) 74.2 20.0 16.7 39.4 55.0 38.6 36.9 57.7 42.3
DeepSeek-Distill-Qwen-7B 90.8 46.7 36.7 56.9 92.5 48.2 48.6 57.1 59.7
+ SPIRAL (Multi-Game) 93.0 43.3 46.7 57.9 92.5 51.1 49.6 58.9 61.7
Multi-game SPIRAL improves Qwen3-4B by 9.2 percentage points (33.1% to 42.3%) and enhances the already-strong DeepSeek-Distill-Qwen-7B model by 2.0 points (59.7% to 61.7%). This demonstrates that game-based reasoning skills provide value even for models that already excel at mathematical reasoning, suggesting that games teach complementary cognitive abilities not captured by traditional training.
Finding 3: Different games develop specialized skills that work together Each game builds distinct abilities that transfer to similar out-of-distribution games. Multi-game training combines these specialized skills, achieving superior performance on novel games that individual specialists struggle with. This synergistic training improves reasoning benchmarks even for already-strong models, demonstrating that games teach complementary cognitive skills.
6.4 RQ4: Is Role-Conditioned Advantage Estimation (RAE) essential for stable self-play training?
We ablate our variance reduction technique to determine whether RAE is necessary to stabilize training and prevent thinking collapse in zero-sum self-play with shared parameters. Figure 6 reveals that without proper variance reduction, self-play training catastrophically fails:
1. Thinking Collapse: The most striking failure occurs in response length (bottom right). Without RAE, reasoning traces plummet from 3,500 to near-zero characters after 200 steps. Models quite literally stop thinking, generating minimal outputs like <think></think><answer>
boxed{bet}</answer>
. This collapse coincides with gradient norms approaching zero, indicating the model has converged to a degenerate policy.
2. Performance Degradation: Without RAE, the model learns solely to play the game while abandoning CoTs, leading to generalization failure. As a result, the math reasoning performance crashes from 35% to 12% at around step 150 (a 66% relative decrease), and the general reasoning drops from 44% to 40%.
3. Training Instability: Policy gradient norms without RAE exhibit higher initial values with spikes, then collapses to near-zero after step 200. In contrast, RAE maintains stable gradient norms of around 0.1 throughout training, enabling continuous learning.
This variance reduction is crucial for maintaining stable reasoning. High-variance gradients cause models to oscillate between extreme strategies, eventually converging to degenerate short responses that minimize variance but abandon reasoning entirely.
Figure 6: Training dynamics comparing REINFORCE with RAE (orange) versus vanilla REINFORCE (gray). RAE maintains stable performance across all metrics while vanilla REINFORCE suffers catastrophic thinking collapse. Top row: Performance on game playing, math reasoning, and general reasoning benchmarks. Bottom row: Policy gradient norm shows training instability without RAE; response length reveals thinking collapse where models stop generating reasoning traces.
Finding 4: RAE prevents models from abandoning reasoning during self-play Without RAE, models began to truncate their reasoning processes after 200 steps, generating empty reasoning traces like <think></think>. This thinking collapse causes math performance to crash while gradients become unstable then drop to zero. RAE maintains stable gradients throughout training, allowing models to keep generating substantive reasoning. Self-play alone is not enough; proper variance reduction is essential for maintaining reasoning in competitive games.
7 Conclusion
Summary of Contributions. We introduced SPIRAL, enabling language models to develop reasoning capabilities through competitive self-play without human-curated data. Our technical contributions include a fully online multi-turn MARL system for LLMs and Role-conditioned Advantage Estimation (RAE), which prevents thinking collapse in zero-sum games. Our empirical results show that training on Kuhn Poker alone improves mathematical reasoning by 8.7% average and Minerva Math by 18.1%, surpassing models trained on 25,000 expert demonstrations. Different games develop distinct transferable skills, and multi-game training achieves superior performance on both reasoning benchmarks and novel out-of-distribution games, demonstrating synergistic benefits from diverse game training.
Limitations. While eliminating human-curated problems, SPIRAL still requires designed game environments. Our experiments use simple games (TicTacToe, Kuhn Poker, Simple Negotiation); scaling to complex environments remains unexplored. The computational requirements are substantial (8 H100 GPUs for 25 hours per experiment). Performance plateaus after extended training, and our evaluation focuses on academic benchmarks rather than real-world reasoning tasks requiring common sense or ethical judgment.
Future Directions. This work opens several avenues: expanding to cooperative games, incorporating partial observability, and designing games targeting specific reasoning weaknesses. Understanding why certain games develop particular skills could enable principled environment design. We envision ecosystems of self-improving agents generating increasingly sophisticated challenges, creating autonomous reasoning development beyond human supervision.
Final Thoughts. SPIRAL demonstrates that simple games can unlock complex reasoning without domain-specific data. By harnessing competitive pressure, we create systems that discover their own curricula and continuously improve. The transfer from gameplay to mathematics could suggest that intelligence emerges not from sophisticated supervision but could from environmental challenges that force models to think. This paradigm shift points toward AI systems that autonomously push reasoning boundaries and continuously evolve through self-play.
Appendix
This appendix provides comprehensive details supporting our main findings. App. A shows how existing SFT and RLVR paradigms adapt to our turn-level MDP formulation. App. B contains implementation specifics including hyperparameters, infrastructure details, and game environment observations referenced in Section 5.2. App. C presents extended results tables and detailed pattern evolution analysis that support our case study in Section 6.1. App. D describes the complete methodology for our reasoning pattern analysis, detailing how we discovered and quantified transfer mechanisms. Finally, App. E provides comprehensive descriptions of for all game environments used in our experiments, including both training and evaluation games.
Appendix A Training Paradigms in Turn-Level MDPs
This section shows how existing training paradigms adapt to our turn-level MDP formulation introduced in Section .
A.1 Supervised Fine-Tuning (SFT) in Turn-Level MDPs
In the turn-level setting, SFT requires a dataset $\mathcal{D}_{\text{SFT}}=\{(s_{i},c_{i}^{*},a_{i}^{*})\}_{i=1}^{N}$ of states with expert reasoning traces $c_{i}^{*}$ and actions $a_{i}^{*}$. The model learns to imitate complete turn-level responses:
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{(s,c^{*},a^{*})\sim\mathcal{D}_{\text{SFT}}}\left[\log\pi_{\theta}(c^{*},a^{*}|s)\right].$ (11)
Note that in single-turn settings where each state $s$ appears only once, SFT reduces to standard behavior cloning. The key limitation remains: SFT requires expensive human annotation of both reasoning traces and final answers.
A.2 Reinforcement Learning with Verifiable Rewards (RLVR) in Turn-Level MDPs
RLVR (DeepSeek Team, 2024) eliminates the need for reasoning supervision, requiring only state-answer pairs $\mathcal{D}_{\text{RLVR}}=\{(s_{i},a_{i}^{*})\}_{i=1}^{N}$. In the turn-level formulation:
$J_{\text{RLVR}}(\theta)=\mathbb{E}_{s\sim\mathcal{D}_{\text{RLVR}},y\sim\pi_{\theta}(\cdot|s)}\left[r(s,a)\right],$ (12)
where $r(s,a)=\mathbb{I}[a=a^{*}]$ indicates answer correctness and $y$ contains both reasoning and action.
In single-turn settings without subsequent interactions, RLVR reduces to a contextual bandit problem (Langford and Zhang, 2007). Recent works on mathematical (Shao et al., 2024; DeepSeek Team, 2024) and code reasoning (Zhu et al., 2024; Xin et al., 2024) show that even this simplified bandit-style RLVR can unlock sophisticated reasoning. However, these approaches still require human-curated problem sets $\mathcal{D}_{\text{RLVR}}$, which SPIRAL eliminates through self-play.
Appendix B Experimental Setup Details
This section provides complete implementation details for reproducing our experiments. We begin with visual examples of game environments, followed by our hyperparameter configurations.
B.1 Game Environment Observations
The language models receive structured text observations from each game environment. Fig. 7 shows example observations from our three training games: TicTacToe, Kuhn Poker, and Simple Negotiation. These observations serve as the input prompts $s_{t}$ at each turn, providing complete game state information in natural language format.
Figure 7: Example observations of three training game environments.
For games with partial observability such as Kuhn Poker and Simple Negotiation, we maintain Markovian state representations by concatenating historical actions into the current state $s_{t}$. This ensures the model has sufficient information for decision-making despite hidden information.
Similarly, Fig. 8 presents observations from five evaluation environments used to test out-of-distribution generalization. These games were never seen during training, allowing us to assess whether learned skills transfer to novel game mechanics.
Figure 8: Example observations of five evaluation game environments.
B.2 Hyperparameter Configuration
Table 7 presents the complete hyperparameter settings used across all experiments. These configurations were selected through preliminary experiments to balance training stability and computational efficiency. Please see our open-source codebase for a complete and reproducible experiment example.
Parameter Value
Actor
Maximum response length $4096$ tokens
Sampling temperature 1.0
(top P, top k) (1.0, -1)
Learner
Optimizer AdamW
Adam parameters ($\beta_{1},\beta_{2}$) (0.9, 0.95)
Weight decay 0.0
Gradient norm clipping 1.0
Batch size 128
Discount factor 1.0
EMA decay rate 0.95
Learning rate scheduler Constant
Learning rate $1\times 10^{-6}$
Inner proximal update epoch 2
KL loss coefficient 0.0
KL penalty coefficient 0.0
Policy clipping parameter 0.2
Table 7: Hyperparameter configurations used in all experiments.
These hyperparameters remain fixed across all game environments and model scales to ensure fair comparison. The distributed training infrastructure utilizes 8 H100 GPUs, with parallel actors generating game trajectories while a centralized learner performs synchronous policy updates.
Appendix C Additional Results and Analysis
C.1 Detailed Evolution of Case-by-Case Analysis
To understand how reasoning patterns develop during training, we tracked the evolution of case-by-case analysis across checkpoints. Table 8 shows a concrete example from Minerva Math Problem 135, illustrating how models progressively develop structured reasoning.
Table 8: Evolution of Case-by-Case Analysis in Minerva Problem 135
Stage Problem: Round cosmic ages to nearest power of 10
Early
(Ckpt 16)
Approach: Lumped reasoning without explicit case separation
[…]
Case Structure: No explicit cases - all subproblems handled in continuous text
Result: Subproblem 2: 14 Gyr $\times$ (should be 10)
Mid
(Ckpt 128)
Approach: Explicit enumeration of cases with structured analysis
<output>
Case 1 - Subproblem 0:
 1. Understanding the Context: BBN at $\sim$20 minutes
 2. Nearest Power of 10: $1010^{1}=$ 10
[…]
Case 2 - Subproblem 1:
 1. Understanding the Context: Recombination at $\sim$377,000 years
 2. Nearest Power of 10: $00010^{5}=100,$ 000
[…]
Case 3 - Subproblem 2:
 1. Understanding the Context: Current age $\sim$13.8 Gyr
 2. Nearest Power of 10: Identifies $10^{10}$ correctly
[…]
Case Structure: Clear case separation with consistent 3-step process
Result: Still incorrect final rounding $\times$
Late
(Ckpt 400)
Approach: Systematic case-by-case with correct execution
<think>
Case 1 - Subproblem 0:
 1. Understand Context: He formation in early universe
 2. Estimate Time: $\sim$20 minutes after Big Bang
[…]
Case 2 - Subproblem 1:
 1. Understand Context: H atom formation (recombination)
 2. Estimate Time: $\sim$380,000 years
[…]
Case 3 - Subproblem 2:
 1. Understand Context: Current universe age
 2. Convert Units: 13.8 billion years = 13.8 Gyr
[…]
Case Structure: Complete systematic enumeration with correct logic
Result: All cases solved correctly $\checkmark$
This progression demonstrates how competitive self-play forces models to develop increasingly structured approaches. Early attempts show unorganized reasoning, while later checkpoints exhibit clear case separation and systematic analysis, a pattern that emerges from game playing and transfers to mathematical problem solving.
C.2 Comprehensive Benchmark Results
Table 9 presents extended results showing SPIRAL’s performance across different training configurations and base models.
Table 9: SPIRAL training improves reasoning benchmarks for different base models
Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average
Qwen3-4B-Base Family
Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 33.1
+ SFT 76.0 16.7 13.3 38.4 50.0 31.2 33.0 48.8 38.4
+ Mistral Opponent (KuhnPoker) 64.0 6.7 6.7 29.8 50.0 26.1 35.6 43.6 32.8
+ Gemini Opponent (KuhnPoker) 69.2 10.0 13.3 33.8 50.0 33.8 35.3 55.5 37.6
+ SPIRAL (TicTacToe) 75.6 10.0 13.3 38.5 55.0 42.6 37.6 57.7 41.3
+ SPIRAL (KuhnPoker) 76.4 13.3 10.0 38.4 57.5 42.4 37.0 56.5 41.4
+ SPIRAL (Negotiation) 76.4 10.0 10.0 41.2 57.5 36.0 34.7 54.2 40.0
+ SPIRAL (TicTacToe+KuhnPoker) 76.2 13.3 16.7 40.7 60.0 41.5 35.7 57.2 42.7
+ SPIRAL (TicTacToe+KuhnPoker+Negotiation) 74.2 20.0 16.7 39.4 55.0 38.6 36.9 57.7 42.3
DeepSeek-Distill-Qwen-7B Family
DeepSeek-Distill-Qwen-7B 90.8 46.7 36.7 56.9 92.5 48.2 48.6 57.1 59.7
+ SPIRAL (Multi-Game) 93.0 43.3 46.7 57.9 92.5 51.1 49.6 58.9 61.7
These results reveal several important insights. First, single-game SPIRAL training (40.0-41.4% average) outperforms supervised fine-tuning on 25,000 expert examples (38.4% average), validating that self-play can discover more effective reasoning strategies than imitating expert demonstrations. Second, multi-game training (42.3-42.7% average) consistently outperforms single-game variants, suggesting that diverse cognitive challenges create more robust reasoning capabilities. Third, SPIRAL improves even strong models like DeepSeek-Distill-Qwen-7B (from 59.7% to 61.7%), demonstrating that competitive game self-play training can enhance models that already excel at reasoning tasks.
Appendix D Case Study Methodology
This section details our systematic approach to discovering and analyzing reasoning pattern transfer from games to mathematics. Rather than searching for predetermined patterns, we employed a bottom-up discovery process to identify what reasoning strategies naturally emerge and transfer between domains.
D.1 Data Collection Framework
Our analysis examined reasoning traces from two sources across three training checkpoints:
Game Trajectories: We collected 290 complete Kuhn Poker games, focusing on winning trajectories to identify successful reasoning strategies. Each trajectory includes the complete thought process from initial card observation through final decision.
Mathematical Solutions: We analyzed 46,792 solution attempts across MATH500, AIME, OlympiadBench, and Minerva Math benchmarks. Solutions were categorized by success (score=1) or failure (score=0) to understand which reasoning approaches prove effective.
Temporal Analysis: Checkpoints at steps 0 (initial), 128 (intermediate), and 400 (final) capture the evolution of reasoning complexity throughout training.
D.2 Bottom-Up Pattern Discovery Process
Rather than searching for predefined patterns, we employed GPT-4.1 to discover patterns that naturally emerge in the data. This bottom-up approach ensures we capture the actual reasoning strategies used rather than imposing our expectations.
Pattern Discovery Prompt Analyze these {domain} reasoning traces using a BOTTOM-UP approach. Don’t look for predefined patterns. Instead, discover what reasoning patterns actually exist. REASONING TRACES ({len(sample)} samples from {len(traces)} total):
{json.dumps([t[’reasoning’] for t in sample[:40]], indent=2)}
Your task: 1. Read through ALL the reasoning traces carefully 2. Identify RECURRING patterns or structures that appear multiple times 3. Group similar reasoning approaches together 4. Name each discovered pattern based on what it actually does 5. Count how many traces use each pattern 6. Provide example quotes for each pattern Format your response as:
PATTERN 1: [Descriptive Name]
- Description: [What this pattern does]
- Count: X/{len(sample)} traces
- Example quotes: [2-3 actual quotes showing this pattern]
Be specific and grounded in the actual data. If you see a pattern only once, don’t include it.
This discovery process revealed three dominant patterns that emerged independently in both domains:
  1. 1.
    Case-by-Case Analysis: Systematic enumeration of scenarios
  2. 2.
    Expected Value Calculation: Probabilistic decision-making
  3. 3.
    Pattern Recognition: Identifying regularities and structures
D.3 Cross-Domain Transfer Quantification
After discovering patterns in each domain, we compared them to identify which strategies transfer between games and mathematics:
Pattern Comparison Prompt Compare the reasoning patterns discovered in games vs mathematics: GAME PATTERNS:
{game_patterns}
MATH PATTERNS:
{math_patterns}
Analyze: 1. Which patterns appear in BOTH domains? (These show transfer) 2. Which patterns are unique to each domain? 3. Calculate transfer rates for shared patterns 4. Identify the most successfully transferred reasoning strategies 5. Explain WHY certain patterns transfer well Focus on concrete evidence of transfer, not speculation.
The transfer analysis revealed:
Case-by-Case Analysis shows near-perfect transfer (72% in games to 71% in math) because systematic enumeration represents domain-agnostic structured thinking. Whether analyzing opponent possibilities in Poker or solution branches in mathematics, the core cognitive skill remains identical.
Expected Value Calculation exhibits limited transfer (78% in games to 28% in math) because explicit probabilistic decision-making appears primarily in probability and optimization problems. Most mathematical domains lack the decision-theoretic structure that makes this pattern universally applicable in games.
Pattern Recognition demonstrates amplification during transfer (35% in games to 45% in math). Mathematics inherently requires pattern identification, so game training enhances an already-essential mathematical skill, producing stronger pattern recognition than games alone develop.
D.4 Pattern Evolution Analysis
To understand how reasoning develops during training, we tracked pattern emergence across checkpoints:
Evolution Analysis Prompt Analyze how reasoning patterns evolve across training checkpoints: {json.dumps(evolution_analysis, indent=2)} For each checkpoint: 1. Describe the complexity of reasoning 2. Identify new patterns that emerge 3. Track how patterns become more sophisticated 4. Show concrete examples of improvement Focus on the actual evolution you can see in the data.
D.5 Concrete Transfer Example Identification
To validate transfer claims, we identified parallel reasoning structures across domains:
Transfer Example Prompt Find the clearest examples of reasoning transfer from games to mathematics: {json.dumps(examples, indent=2)} For each complexity level (short/medium/long): 1. Identify parallel reasoning structures 2. Quote specific passages that show transfer 3. Explain what cognitive skill is being transferred 4. Rate the clarity of transfer (1-10) Focus on examples where the same reasoning approach clearly appears in both domains.
D.6 Pattern Classification at Scale
After discovering patterns through bottom-up analysis, we classified all traces to measure transfer rates:
Pattern Classification Prompt Analyze these {domain} reasoning traces and find examples of each pattern. PATTERNS TO FIND: 1. Case-by-Case Analysis: Systematic enumeration of different scenarios/cases 2. Expected Value Calculation: Explicit probability calculations, computing expected outcomes 3. Pattern Recognition: Identifying recurring structures, noticing trends REASONING TRACES:
{json.dumps(batch, indent=2)}
For EACH pattern, identify ALL examples that clearly demonstrate it.
Return in this EXACT format:
CASE_BY_CASE_INDICES: [list of indices]
EXPECTED_VALUE_INDICES: [list of indices]
PATTERN_RECOGNITION_INDICES: [list of indices]
Be strict - only include clear examples of each pattern.
D.7 Validation Methodology
To ensure robust findings, we implemented multiple validation steps:
Sampling Strategy: We analyzed 50 random trajectory samples per checkpoint to avoid selection bias while maintaining computational feasibility.
Success Stratification: Separate analysis of successful and failed attempts revealed which reasoning strategies genuinely contribute to problem-solving rather than merely appearing frequently.
Manual Verification: Spot-checking GPT-4.1’s pattern classifications against raw traces confirmed the accuracy of automated analysis.
Scale Validation: After discovering patterns through focused analysis, we classified all 46,792 mathematical traces to verify that observed transfer rates hold at scale.
This methodology ensures our findings reflect genuine cognitive transfer rather than superficial pattern matching, providing quantitative evidence that competitive gameplay develops reasoning skills applicable far beyond the training domain.
Appendix E Game Environment Specifications
This section provides detailed specifications for all game environments used in our experiments, including both training and evaluation games.
E.1 Training Game Environments
TicTacToe tests spatial pattern recognition through perfect information gameplay. Players alternate placing marks on a 3×3 grid, aiming to create lines of three. The deterministic nature isolates pure strategic reasoning from uncertainty management. Success requires recognizing winning patterns, blocking opponent threats, and creating fork positions that guarantee victory. We hypothesize these skills transfer to geometric reasoning and spatial visualization tasks.
Kuhn Poker introduces probabilistic reasoning through minimal hidden information. With only three cards (Jack, Queen, King), one per player plus one undealt, the game distills poker to essential elements of bluffing and value betting. Players can check, bet, call, or fold, with outcomes determined by card strength. Success requires calculating expected values, modeling opponent behavior, and making decisions under uncertainty. These capabilities should transfer to probability problems and strategic decision-making.
Simple Negotiation develops multi-constraint optimization through resource trading. Two players exchange Wood and Gold tokens, with opposing utility functions creating natural tension. Each player aims to maximize portfolio value through proposals and counteroffers. Success requires understanding opponent preferences, planning multi-step trades, and strategic communication. We expect these skills to enhance optimization problems and multi-constraint reasoning tasks.
E.2 Out-of-Distribution Evaluation Games
Our evaluation suite tests whether learned skills generalize to novel mechanics:
Snake extends spatial reasoning to dynamic environments. Players control snakes navigating grids to collect apples while avoiding collisions with walls, themselves, or opponents. This tests whether static pattern recognition from TicTacToe transfers to trajectory planning and dynamic obstacle avoidance.
Connect Four adds directional constraints to spatial reasoning. Players drop discs into columns with gravity, aiming to connect four vertically, horizontally, or diagonally. The vertical placement mechanic tests whether TicTacToe strategies generalize to games with restricted movement options.
Pig Dice isolates risk-reward decision making. Players repeatedly roll dice to accumulate points but lose all turn points when rolling 1. This tests whether probabilistic reasoning from Kuhn Poker extends to sequential risk assessment and expected value calculation in different contexts.
Liar’s Dice combines probability with deception detection. Players make escalating claims about hidden dice totals and challenge suspected bluffs. Success requires probability calculation, opponent modeling, and strategic deception, testing whether Kuhn Poker skills scale to multi-round bluffing games.
Truth and Deception focuses on asymmetric information and persuasion. One player knows the true fact among options and misleads through conversation while the other must identify truth through questioning. This evaluates whether negotiation skills transfer to pure communication strategy.
These diverse evaluation games probe different aspects of transfer learning, revealing which cognitive skills generalize beyond their training context and confirming that SPIRAL develops fundamental reasoning capabilities rather than game-specific tactics.

### v3

@@
-SPIRAL: Self-Play on Zero-Sum Games Incentivizes Reasoning via Multi-Agent Multi-Turn Reinforcement Learning
+arXiv:2506.24119v3[cs.AI] 02 Mar 2026
 Abstract
 Recent advances in reinforcement learning have shown that language models can develop sophisticated reasoning through training on tasks with verifiable rewards, but these approaches depend on human-curated problem-answer pairs and domain-specific reward engineering.
-We introduce SPIRAL, a self-play framework where models learn by playing multi-turn, zero-sum games against continuously improving versions of themselves, eliminating the need for human supervision.
-Through self-play, SPIRAL generates an infinite curriculum of progressively challenging problems as models must constantly adapt to stronger opponents.
+We introduce SPIRAL, a self-play framework where models learn by playing multi-turn, zero-sum games against continuously improving versions of themselves, generating an automatic curriculum of stronger opponents, and eliminating the need for human supervision.
 To enable this self-play training at scale, we implement a fully online, multi-turn, multi-agent reinforcement learning system for LLMs and propose role-conditioned advantage estimation (RAE) to stabilize multi-agent training.
-Using SPIRAL, self-play on zero-sum games produces reasoning capabilities that transfer broadly.
-Training Qwen3-4B-Base on Kuhn Poker alone achieves 8.6% improvement on math and 8.4% on general reasoning, outperforming SFT on 25,000 expert game trajectories.
-Analysis reveals that this transfer occurs through three cognitive patterns: systematic decomposition, expected value calculation, and case-by-case analysis.
-Multi-game training (TicTacToe, Kuhn Poker, Simple Negotiation) further enhances performance as each game develops distinct reasoning strengths.
-Applying SPIRAL to a strong reasoning model (DeepSeek-R1-Distill-Qwen-7B) can still lead to 2.0% average improvement.
-These results demonstrate that zero-sum games naturally develop transferable reasoning capabilities, highlighting a promising direction for autonomous reasoning development.
-[Uncaptioned image] Spiral-RL/SPIRAL       [Uncaptioned image] Spiral-RL
+SPIRAL produces reasoning capabilities that transfer broadly, improving performance by up to 10% across a suite of 8 reasoning benchmarks on 4 different models spanning Qwen and Llama model families, outperforming supervised fine-tuning on 25,000 expert game trajectories.
+Multi-game training (TicTacToe, Kuhn Poker, Simple Negotiation) yields the strongest results, with improvements observed across both base and instruction-tuned models.
+Analysis of chain-of-thought traces reveals that games develop distinct cognitive patterns that transfer to improve reasoning performance, with different games developing complementary strengths.
+Even models which have already been trained on reasoning tasks using RLVR, like DeepSeek-R1-Distill-Qwen-7B, still benefit from our approach.
+These results demonstrate that zero-sum games naturally develop transferable reasoning capabilities across diverse model architectures and training stages, highlighting a promising direction for autonomous reasoning development.
+Our code can be found in https://github.com/spiral-rl/spiral.
 Figure 1: SPIRAL achieves consistent improvements over base models across game performance and reasoning benchmarks.
@@
 Through outcome-based rewards, RL enables models to develop generalizable reasoning strategies and consistently solve complex problems where supervised fine-tuning shows limited progress.
-However, current approaches to reasoning enhancement face a fundamental scalability bottleneck: their dependence on carefully engineered reward functions, domain-specific datasets, and expert supervision (DeepSeek Team, 2024; Ouyang et al., 2022; Bai et al., 2022).
+However, current approaches face a fundamental scalability bottleneck: dependence on carefully engineered reward functions, domain-specific datasets, and expert supervision (DeepSeek Team, 2024; Ouyang et al., 2022; Bai et al., 2022).
 Each new reasoning domain requires experts to craft evaluation metrics, curate training problems, and validate reasoning traces.
-This manual process becomes increasingly unsustainable as we pursue more general intelligence, limiting both the scale and diversity of reasoning challenges that models can learn from.
-Self-play offers a solution to this scalability crisis by eliminating the need for human supervision in training data creation (Silver et al., 2017; Tesauro, 1995).
-In self-play, models learn by competing against copies of themselves, where each match outcome provides automatic feedback.
-As the model improves, its opponent improves equally, maintaining a consistent challenge level that drives continuous learning.
-This paradigm has already revolutionized AI across many domains: from TD-Gammon’s backgammon supremacy (Tesauro, 1995) to AlphaGo’s conquest of Go (Silver et al., 2016, 2017) to OpenAI Five’s mastery of complex team coordination (Berner et al., 2019).
-Yet despite these successes, applying self-play to enhance language model reasoning remains largely unexplored.
-Prior attempts have been limited to simple word games with offline updates (Cheng et al., 2024), LoRA adaptations that constrain learning capacity (Dettmers et al., 2023; Park et al., 2025), or single-turn code generation tasks (Zhao et al., 2025), falling short of leveraging multi-turn competitive dynamics that enable extended strategic reasoning.
-We introduce SPIRAL, which applies self-play to two-player zero-sum language games for developing reasoning capabilities.
-Zero-sum language games provide ideal training environments: they require strategic thinking and planning while remaining simple enough to enable stable learning, with clear rules that make outcomes verifiable.
-SPIRAL offers two key advantages.
-First, unlike traditional RLVR approaches that depend on human-curated problem-answer pairs, it generates infinite training data through game dynamics alone.
-Second, compared to fixed-opponent training (see Fig.
-2), self-play prevents overfitting to static strategies by continuously evolving the challenge level.
-However, implementing this vision for LLMs presents significant technical challenges.
+This manual process becomes increasingly unsustainable as we pursue more general intelligence, limiting both scale and diversity of reasoning challenges that models can learn from.
+Self-play on games offers a solution by eliminating human supervision in training data creation (Silver et al., 2017; Tesauro, 1995).
+In game-based self-play, models learn by competing against copies of themselves, where game outcomes provide automatic feedback and opponents improve equally, maintaining a consistent challenge that drives continuous learning.
+Although many prominent successes in past AI research relied on self-play—from TD-Gammon (Tesauro, 1995) to AlphaGo (Silver et al., 2016; 2017) to OpenAI Five (Berner et al., 2019)—so far, applying self-play on games to enhance language model reasoning remains largely unexplored.
+Prior attempts have been limited to simple word games with offline updates (Cheng et al., 2024), LoRA adaptations (Dettmers et al., 2023; Park et al., 2025), or single-turn tasks (Zhao et al., 2025), falling short of leveraging multi-turn competitive dynamics for extended strategic reasoning.
+We introduce SPIRAL (Self-Play on zero-sum games Incentivizes Reasoning via multi-Agent multi-turn reinforcement Learning), which applies self-play to two-player zero-sum language games for developing reasoning capabilities.
+SPIRAL offers two key advantages: unlike traditional RLVR approaches depending on human-curated problem-answer pairs, it generates unlimited training data through game dynamics alone; compared to fixed-opponent training (see Fig.˜2), self-play prevents overfitting to static strategies by continuously evolving challenge level.
+However, implementing this for LLMs presents significant challenges.
 The computational demands of multi-turn, multi-agent autoregressive generation require sophisticated distributed systems, while standard RL algorithms suffer from high variance in multi-agent settings.
-To address these challenges, we implement a fully online, multi-turn, multi-agent reinforcement learning system with a distributed actor-learner architecture.
-We also introduce role-conditioned advantage estimation (RAE), which stabilizes training by normalizing rewards relative to each player’s expected performance.
-Without RAE, our ablation experiments show that models suffer from thinking collapse, progressively abandoning reasoning traces that are critical for generalization.
+We address these through a fully online, multi-turn, multi-agent reinforcement learning system with distributed actor-learner architecture and introduce role-conditioned advantage estimation (RAE), which stabilizes training by normalizing rewards relative to each player’s expected performance.
 Key Findings.
-Training on zero-sum games produces reasoning capabilities that transfer broadly.
-SPIRAL trained solely on Kuhn Poker improves Qwen3-4B-Base by 8.6% on mathematical reasoning and 8.4% on general reasoning, outperforming SFT on 25,000 expert game trajectories despite using only game states with no mathematical content.
-This transfer works through three cognitive patterns we identify: systematic decomposition, expected value calculation, and case-by-case analysis.
-These patterns develop effectively through self-play, as fixed-opponent training fails while self-play continuously improves via adaptive challenge.
-Building on this, different games develop complementary skills (TicTacToe for spatial reasoning, Kuhn Poker for probabilistic thinking, and Simple Negotiation for strategic optimization) that combine synergistically in multi-game training.
-All these benefits require our Role-conditioned Advantage Estimation (RAE); without it, models abandon reasoning after 200 steps.
-The approach generalizes beyond weak models, with SPIRAL improving even strong models like DeepSeek-R1-Distill-Qwen-7B by 2.0%.
-These results demonstrate that zero-sum games can serve as powerful reasoning gymnasiums, developing cognitive skills that transfer far beyond their training domain.
+Training on zero-sum games produces reasoning capabilities that transfer broadly across diverse model architectures.
+Multi-game SPIRAL training (TicTacToe, Kuhn Poker, Simple Negotiation) achieves up to 10% improvement across 8 reasoning benchmarks, outperforming supervised fine-tuning on 25,000 expert trajectories.
+On Qwen3-4B-Base (Yang et al., 2025), multi-game training reaches 44.5% average performance versus 34.0% baseline (+10.5+10.5% absolute gain), while Qwen3-8B-Base (Yang et al., 2025) improves from 39.5% to 49.6% (+10.1+10.1%).
+The approach generalizes across model families: base models (Qwen3-4B/8B-Base, Octothinker-8B-Base; Wang et al.
+(2025a)) and instruction-tuned models (Llama-3.1-8B-Instruct; Dubey et al.
+(2024)) all show consistent improvements, with Octothinker-8B-Base gaining 8.0% and Llama-3.1-8B-Instruct improving 2.0% despite already being instruction-tuned.
+Each game develops complementary cognitive skills: TicTacToe for spatial reasoning, Kuhn Poker for probabilistic thinking, and Simple Negotiation for strategic optimization, which combine synergistically in multi-game training.
+Using post-hoc analysis, we find examples of three patterns learned from gameplay that transfer to improve math performance: case-by-case analysis, expected value calculation, and pattern recognition.
+These patterns develop effectively through self-play’s adaptive curriculum, as fixed-opponent training fails while self-play continuously improves.
+Role-conditioned Advantage Estimation proves critical: without RAE, models abandon reasoning after 200 steps, progressively generating empty thinking traces that destroy generalization.
 Building on these findings, our work makes the following contributions:
+  1. 1.
+    A Fully Online, Multi-Turn, Multi-Agent RL Framework for LLMs: We develop a distributed actor-learner architecture enabling online self-play with full-parameter updates across multiple two-player zero-sum language games.
+The multi-turn aspect trains models to reason through sequential decisions, directly preparing them for complex multi-step problem solving.
+Unlike prior offline approaches, this provides continuous curriculum as the model adapts to an ever-improving opponent.
+We release our implementation to facilitate further research.
+  2. 2.
+    Role-conditioned Advantage Estimation (RAE): We introduce a variance-reduced advantage estimator specifically designed for multi-agent settings.
+By normalizing rewards relative to each player’s expected performance, RAE prevents the degradation of the model’s reasoning capabilities, a failure mode we term “thinking collapse”.
+Without it, models progressively abandon reasoning traces after 200 steps, which is critical for generalization.
+  3. 3.
+    Empirical Discovery of Transfer: We demonstrate that self-play on zero-sum games improves both out-of-distribution game performance and academic reasoning benchmarks by up to 10% without domain-specific training data.
+Our analysis identifies reasoning patterns (systematic decomposition, expected value calculation, case-by-case analysis) that transfer from games to mathematics at measurable rates, with different games developing specialized skills that combine synergistically in multi-game training.
 Figure 2: From human-designed rewards to self-discovered reasoning through SPIRAL.
@@
 Right: SPIRAL enables continuous reasoning improvement through self-play, where both players develop increasingly sophisticated strategies without human supervision.
-  1. 1.
-    A Fully Online, Multi-Turn, Multi-Agent RL Framework for LLMs: We develop a distributed actor-learner architecture that enables online self-play with full-parameter updates across multiple two-player zero-sum language games.
-The multi-turn aspect trains models to reason through sequential decisions, directly preparing them for complex multi-step problem solving.
-Unlike prior offline approaches, this provides a continuous curriculum as the model must adapt to an ever-improving opponent: itself.
-We release our implementation111 https://github.com/spiral-rl/spiral.
-to facilitate further research.
-  2. 2.
-    Role-conditioned Advantage Estimation (RAE): We introduce a variance-reduced advantage estimator specifically designed for multi-agent settings.
-By normalizing rewards relative to each player’s expected performance, RAE reduces variance in multi-agent learning.
-Our ablation studies empirically validate that RAE stabilizes learning and prevents thinking collapse.
-Without it, models progressively abandon reasoning traces, which is critical for generalization.
-  3. 3.
-    Empirical Discovery of Transfer: We demonstrate that self-play on zero-sum games improves both out-of-distribution game performance and academic reasoning benchmarks without domain-specific training data.
-Our analysis identifies reasoning patterns (systematic decomposition, expected value calculation, case-by-case analysis) that transfer from games to mathematics at measurable rates.
-Different games develop specialized skills that transfer to related domains, and multi-game training combines these abilities synergistically.
 2 Related Work
 Reinforcement Learning for LLM Reasoning.
-Reinforcement learning has evolved from alignment tasks using RLHF (Jaques et al., 2019; Ouyang et al., 2022; Bai et al., 2022) to directly improving reasoning capabilities.
+Reinforcement learning (RL) in LLMs has progressed from alignment tasks using RLHF (Jaques et al., 2019; Ouyang et al., 2022; Bai et al., 2022) to directly improving reasoning capabilities.
 Recent models like OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek Team, 2024) demonstrate that RL with verifiable rewards (RLVR) can unlock chain-of-thought reasoning using rule-based rewards (Lightman et al., 2023; Uesato et al., 2022).
 However, these approaches depend on human-curated problem sets and domain-specific reward engineering.
-SPIRAL eliminates this dependency by using self-play games to generate infinite reasoning challenges without human supervision.
-Multi-Agent RL for Language Models.
-Implementing MARL for full-scale LLMs presents significant technical challenges (Wan et al., 2025; Liu et al., 2025a).
-Prior work has made various compromises: Sarkar et al.
-(2025) uses RNNs instead of transformers; Jacob et al.
-(2022) focuses on simplified environments that don’t require full autoregressive generation; and Liao et al.
-(2024) shows the effectiveness of self-play with SFT on proprietary models.
+SPIRAL eliminates this dependency by using self-play games to generate unlimited reasoning challenges without human supervision.
+Self-Play and Multi-Agent RL for LLMs.
+Self-play in LLMs initially focused on alignment objectives (Chen et al., 2024; Yuan et al., 2024) before recent work applied it to enhance model capabilities.
+SPAG (Cheng et al., 2024) applies self-play to Adversarial Taboo using offline updates on a single game; SPC (Chen et al., 2025) and Genius (Xu et al., 2025a) require predefined human task distributions; Absolute Zero (Zhao et al., 2025) generates single-turn coding tasks; Foundation Model Self-Play (Dharna et al., 2025) uses foundation models to evolve code-based policies rather than direct gameplay; Prover-Verifier Game (Kirchner et al., 2024) improve output legibility through adversarial training.
+Implementing multi-agent RL (MARL) for full-scale LLMs presents significant technical challenges (Wan et al., 2025; Liu et al., 2025b; a).
+Prior work circumvents these challenges by using RNNs instead of transformers (Sarkar et al., 2025), restricting to simplified environments (Jacob et al., 2022; Sukhbaatar et al., 2018), or applying supervised fine-tuning on trajectories from proprietary models (Liao et al., 2024).
 In contrast, SPIRAL implements fully online, full-parameter MARL through a distributed actor-learner architecture, enabling continuous adaptation to evolving opponents across multiple games.
-Self-Play for Autonomous Improvement.
-Self-play in LLMs began with alignment methods like SPIN (Chen et al., 2024) and Self-Rewarding Language Models (Yuan et al., 2024).
-Recent work explores capability improvement: SPAG (Cheng et al., 2024) applies self-play to Adversarial Taboo but uses offline updates and remains confined to a single word game; SPC (Chen et al., 2025) and Genius (Xu et al., 2025) require human task distributions; Absolute Zero (Zhao et al., 2025) generates coding tasks but needs deterministic verification.
-SPIRAL demonstrates that self-play on strategic games improves academic benchmarks by 8.7% average without domain-specific data.
 LLMs in Gaming.
-Games serve as testbeds for evaluating LLM capabilities (Paglieri et al., 2024; Ruoss et al., 2024; Zhang et al., 2024; Duan et al., 2024), but recent work explores games as training domains (Feng et al., 2024).
-LMRL-Gym (Abdulhai et al., 2023) benchmarks RL algorithms for LLMs in game environments; RAGEN (Wang et al., 2025) uses single-agent multi-turn RL; ViGaL (Xie et al., 2025) shows arcade games improve math reasoning; Divide-Fuse-Conquer (Zhang et al., 2025) trains on grouped games using offline learning.
-SPIRAL uniquely combines: (1) multi-agent self-play where both players share parameters, (2) fully online learning with continuous opponent evolution, and (3) transfer from zero-sum language games to reasoning without seeing any benchmark-related problems during training.
-3 Preliminaries
-Turn-level Markov Decision Process (MDP).
-Language model training traditionally formulates generation as a token-level MDP (Bellman, 1957; Rafailov et al., 2024) where each action is a single token from vocabulary $\mathcal{V}$.
-For multi-turn reasoning and game-playing, we instead adopt a turn-level MDP formulation $\mathcal{M}=(\mathcal{S},\mathcal{A},T,r,\gamma)$.
-Here, states $\mathcal{S}$ represent complete contexts (e.g., game configurations, problem states, or conversation histories), actions $\mathcal{A}$ are complete responses (containing many tokens), the transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ determines state dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides immediate rewards, and $\gamma\in[0,1]$ is the discount factor.
-The return is defined as the discounted sum of rewards: $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r_{t}$.
-The key distinction: in token-level MDPs, each decision outputs one token; in turn-level MDPs, each decision produces a complete multi-token response before transitioning.
-At each turn $t$, the language model observes state $s_{t}$ and generates:
-$y_{t}=\langle\text{think}\rangle c_{t}\langle/\text{think}\rangle\langle\text{answer}\rangle a_{t}\langle/\text{answer}\rangle,$ (1)
-where $c_{t}$ externalizes reasoning and $a_{t}\in\mathcal{A}$ is the executable action. (See App. A for how existing SFT and RLVR paradigms adapt to turn-level MDPs.)
-Two-Player Zero-Sum Markov Games.
-We extend the single-agent MDP to competitive settings with a two-player zero-sum Markov game (Littman, 1994) $\mathcal{G}=(\mathcal{S},\mathcal{A}_{0},\mathcal{A}_{1},T,r,\gamma)$, where $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are the action spaces for player 0 and player 1 respectively.
-The zero-sum property requires:
-$r_{0}(s,a^{(0)},a^{(1)})+r_{1}(s,a^{(0)},a^{(1)})=0\quad\forall s,a^{(0)},a^{(1)},$ (2)
-where $a^{(0)}\in\mathcal{A}_{0}$ and $a^{(1)}\in\mathcal{A}_{1}$ denote actions taken by each player.
-Given trajectory $\tau=\{(s_{t},a_{t}^{(0)},a_{t}^{(1)})\}_{t=0}^{T}$, the returns satisfy $R_{1}(\tau)=-R_{0}(\tau)$.
+Games serve as both evaluation benchmarks (Paglieri et al., 2024; Ruoss et al., 2024; Zhang et al., 2024; Duan et al., 2024) and training domains (Feng et al., 2024; Verma et al., 2025).
+LMRL-Gym (Abdulhai et al., 2023) and RAGEN (Wang et al., 2025b) both employ single-agent multi-turn RL, with LMRL-Gym providing 8 benchmarking tasks and RAGEN focusing on trajectory-level optimization.
+ViGaL (Xie et al., 2025b) shows that single-agent RL on visual-spatial games transfers to mathematical reasoning without explicit math training.
+Logic-RL (Xie et al., 2025a) trains on synthetic puzzle games; Divide-Fuse-Conquer (Zhang et al., 2025) applies offline learning to grouped games; Boundless Socratic Learning (Schaul, 2024) uses language games for continual learning; Code2Logic (Tong et al., 2025) synthesizes reasoning data from game code.
+We also distinguish SPIRAL from works designed for achieving super-human performance at specific games.
+Cicero (FAIR et al., 2022) integrates a language model with a separate strategic planning algorithm to achieve human-level performance in Diplomacy.
+Similarly, agents developed for Werewolf (Xu et al., 2024; 2025b) and Avalon (Wang et al., 2023) focus on optimizing strategic communication and hidden-role deduction to win within their respective game rules.
+While these works target in-domain victory, SPIRAL treats competitive pressure as a training scaffold to develop reasoning patterns (e.g., case analysis) that transfer to out-of-domain reasoning tasks.
+SPIRAL uniquely combines three elements: (1) multi-agent self-play where both players share parameters, (2) fully online learning with continuous opponent evolution, and (3) demonstrated transfer from zero-sum language games to academic reasoning benchmarks achieving up to 10.5% improvement without exposure to benchmark-related problems during training.
 3 The SPIRAL Framework
@@
 These are then used for on-policy reinforcement learning updates.
-We present SPIRAL, a framework that enables language models to develop generalizable reasoning capabilities through multi-turn competitive self-play. The framework is illustrated in Fig. 3.
-4.1 SPIRAL as Self-Play on Turn-Based Zero-Sum Games
-SPIRAL implements self-play through turn-based zero-sum language games from a collection $\mathcal{G}=\{G_{1},G_{2},...,G_{n}\}$.
-Each game $G_{i}$ is a two-player zero-sum Markov game, with the specific structure that players alternate turns rather than acting simultaneously.
-SPIRAL leverages three properties of these games:
-First, the multi-turn structure mirrors sequential reasoning problems.
-Players alternate turns: at time $t$, player $p=t\bmod 2$ acts while the opponent waits, formally: $a_{t}^{(p)}\in\mathcal{A}_{p}$ and $a_{t}^{(1-p)}=\emptyset$.
-Unlike single-turn bandit settings, this framework trains models to maintain context, plan ahead, and adapt strategies based on evolving game states.
-Second, the zero-sum competitive dynamics create pressure for continuous improvement.
-Each game uses sparse terminal rewards: $r_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})=0$ for all non-terminal states.
-At terminal state $s_{T}$, players receive opposite rewards:
-$R_{0}(\tau)=\rho_{i}(s_{T}),\quad R_{1}(\tau)=-\rho_{i}(s_{T}),$ (3)
-where $\rho_{i}:\mathcal{S}_{i}^{\text{terminal}}\rightarrow\{-1,0,1\}$ determines the outcome. This opposition forces models to develop robust strategies that cannot be easily exploited.
-Finally, both players share a single policy $\pi_{\theta}$ with role conditioning.
-To enable this in the LLM setting, we prepend role-specific system prompts that specify whether the model is playing as Player 0 or Player 1 (see Sec.
-B.1 for example observations).
-At each turn, the active player generates:
-$y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i}),$ (4)
-where conditioning on $p$ is achieved through the system prompt [^2] We note that both $p$ and $G_{i}$ can be subsumed into state $s_{t}$, but we choose to make them explicit here for clarity..
-From $y_{t}^{(p)}$, we extract the action $a_{t}^{(p)}$ to update the game state.
-This shared-parameter approach ensures that as the model improves at one role, it simultaneously faces a stronger opponent, creating an automatic curriculum.
-4.2 Multi-Turn Self-Play Training Objective
-In self-play, we could train two separate policies $\pi_{\theta_{0}}$ and $\pi_{\theta_{1}}$ for each player. Each would maximize their own expected return:
-$J_{0}(\theta_{0})$ $=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta_{0}}\times\pi_{\theta_{1}}|G}\left[R_{0}(\tau)\right]$ (5)
-$J_{1}(\theta_{1})$ $=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta_{0}}\times\pi_{\theta_{1}}|G}\left[R_{1}(\tau)\right]$ (6)
-However, SPIRAL uses a single shared policy $\pi_{\theta}$ for both players, setting $\theta_{0}=\theta_{1}=\theta$.
-While this creates opposing objectives since $R_{1}(\tau)=-R_{0}(\tau)$ in zero-sum games, the role conditioning through system prompts enables the model to learn distinct strategies for each position within a unified policy framework.
-To optimize this shared policy, we apply policy gradient methods. Using REINFORCE (Williams, 1992), the gradient becomes:
-$\nabla_{\theta}J(\theta)=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta}\times\pi_{\theta}|G}\left[\underbrace{\sum_{t\in T_{0}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(0)}|s_{t},0,G)\cdot R_{0}(\tau)}_{\text{Player 0 gradient}}+\underbrace{\sum_{t\in T_{1}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(1)}|s_{t},1,G)\cdot R_{1}(\tau)}_{\text{Player 1 gradient}}\right],$ (7)
+We present SPIRAL, a framework enabling language models to develop generalizable reasoning through multi-turn competitive self-play on games, illustrated in Figure 3.
+Formulation.
+SPIRAL implements self-play through turn-based zero-sum language games from collection $\mathcal{G}=\{G_{1},G_{2},...,G_{n}\}$.
+Each game $G_{i}$ is a two-player zero-sum Markov game (Littman, 1994) built on turn-level MDPs where states $s\in\mathcal{S}$ represent complete contexts (e.g., game configurations), actions $a\in\mathcal{A}$ are complete multi-token responses, and transition function $T_{i}$ determines state dynamics after full turn completion.
+The zero-sum property ensures $r_{0}(s,a^{(0)},a^{(1)})+r_{1}(s,a^{(0)},a^{(1)})=0$ for all states and actions where $a^{(p)}$ denotes the action of player $p\in\{0,1\}$, creating competitive dynamics.
+See Appendix C for detailed formulations.
+Benefits of multi-turn, zero-sum games.
+Zero-sum dynamics create continuous improvement pressure through rewards given only at game termination: $r_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})=0$ for all non-terminal states, with terminal rewards $R_{0}(\tau)=\rho_{i}(s_{T})$ and $R_{1}(\tau)=-\rho_{i}(s_{T})$ where $\rho_{i}:\mathcal{S}_{i}^{\text{terminal}}\rightarrow\{-1,0,1\}$ determines the outcome and $\tau$ represents the complete trajectory.
+This forces robust strategy development as models only receive feedback upon game completion.
+The multi-turn structure mirrors sequential reasoning problems: players alternate turns with $p=t\bmod 2$ acting at time $t$ while the opponent waits, training models to maintain context, plan ahead, and adapt strategies.
+Self-play.
+Rather than training separate policies $\pi_{\theta_{0}}$ and $\pi_{\theta_{1}}$ for each player, SPIRAL uses a single shared policy $\pi_{\theta}$ with parameters $\theta$, setting $\theta_{0}=\theta_{1}=\theta$.
+Role conditioning through system prompts enables the model to learn distinct strategies for each position (see Appendix D.1).
+At each turn, the active player generates a complete response $y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i})$ conditioned on current state $s_{t}$, player role $p$, and game $G_{i}$.
+From this response, we extract the action $a_{t}^{(p)}$ to update the game state via $s_{t+1}=T_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})$ where $a_{t}^{(1-p)}=\emptyset$ for the inactive player.
+This shared-parameter approach ensures efficient use of GPU memory while also guaranteeing that as the model improves at one role, it simultaneously faces a stronger opponent, creating an automatic curriculum.
+Algorithm 1 presents the complete training procedure.
+RL objective. To optimize this shared policy, we apply Monte Carlo policy gradient methods. Using REINFORCE (Williams, 1992), the gradient becomes:
+$\nabla_{\theta}J(\theta)=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta}\times\pi_{\theta}|G}\left[\sum_{t\in T_{0}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(0)}|s_{t},0,G)\cdot R_{0}(\tau)+\sum_{t\in T_{1}}\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(1)}|s_{t},1,G)\cdot R_{1}(\tau)\right],$ (1)
 where $T_{p}=\{t:t\bmod 2=p\}$ denotes turns where player $p$ acted.
-This formulation uses Monte Carlo returns which suffer from high variance, particularly problematic in self-play where the opponent’s strategy continuously evolves, making the environment highly non-stationary.
+This formulation uses Monte Carlo returns which suffer from high variance, particularly problematic in self-play where the opponent’s strategy continuously evolves, making the environment non-stationary.
 Algorithm 1 SPIRAL: Role-Balanced Multi-Turn Self-Play
@@
 8:   for turn $t=0,1,2,$... until terminal do
-9:              $p\leftarrow t\bmod 2$triangleright Determine active player
-10:              $y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i})$triangleright Generate reasoning + action
+9:     $p\leftarrow t\bmod 2$ $\triangleright$ Determine active player
+10:     $y_{t}^{(p)}\sim\pi_{\theta}(\cdot|s_{t},p,G_{i})$ $\triangleright$ Generate ‘[reasoning] [action]
 11:     $a_{t}^{(p)}\leftarrow\text{extract\_action}(y_{t}^{(p)})$
-12:              $a_{t}^{(1-p)}\leftarrow\emptyset$triangleright Inactive player
+12:     $a_{t}^{(1-p)}\leftarrow\emptyset$ $\triangleright$ Inactive player
 13:     $s_{t+1}\leftarrow T_{i}(s_{t},a_{t}^{(0)},a_{t}^{(1)})$
@@
 15:   $R_{0}\leftarrow\rho_{i}(s_{T})$, $R_{1}\leftarrow-R_{0}$
-16:         Add $(\tau,G_{i})$ to batch $\mathcal{B}$triangleright Store trajectory with its game
-17:     end for
-18:     // Role-Balanced Policy Optimization
-19:     for $(\tau,G_{i})\in\mathcal{B}$ do
-20:         for $p\in\{0,1\}$ do
-21:              $b_{G_{i},p}\leftarrow\alpha b_{G_{i},p}+(1-\alpha)R_{p}(\tau)$
-22:              $A_{G_{i},p}(\tau)\leftarrow R_{p}(\tau)-b_{G_{i},p}$
-23:         end for
+16:   Define $\tau=\{(s_{0},y_{0}),(s_{1},y_{1}),\dots,(s_{T},y_{T})\}$
+17:   Add $(\tau,G_{i})$ to batch $\mathcal{B}$ $\triangleright$ Store trajectory with its game
+18:  end for
+19:  // Role-Balanced Policy Optimization
+20:  for $(\tau,G_{i})\in\mathcal{B}$ do
+21:   for $p\in\{0,1\}$ do
+22:     $b_{G_{i},p}\leftarrow\alpha b_{G_{i},p}+(1-\alpha)R_{p}(\tau)$
+23:     $A_{G_{i},p}(\tau)\leftarrow R_{p}(\tau)-b_{G_{i},p}$
 24:   end for
-25:     Update $\theta$ using policy gradient with advantages $A_{G_{i},p}$ (Eq. 10)
-26:end while
-4.3 Policy Optimization with Role-conditioned Advantage Estimation
+25:  end for
+26:  Update $\theta$ on full sequences $y_{t}$ using REINFORCE with advantages $A_{G_{i},p}$ (Eq. 3)
+27:end while
+Role-conditioned advantage estimation.
+Self-play on zero-sum games implies using the same model to optimize for opposing objectives, since $R_{1}(\tau)=-R_{0}(\tau)$.
+This can lead to unstable training dynamics which impedes learning.
 To reduce the high variance inherent in multi-agent REINFORCE, we introduce Role-conditioned Advantage Estimation (RAE).
 In two-player games, even with a shared policy, different roles may have different expected returns due to game asymmetries (e.g., first-move advantage in TicTacToe, information asymmetry in Kuhn Poker).
-RAE maintains separate baselines $b_{G,p}$ for each game $G\in\mathcal{G}$ and role $p\in\{0,1\}$, where each baseline estimates the expected return $\mathbb{E}[R_{p}(\tau)]$ for that role in that game.
-We update these baselines using exponential moving average (EMA) with a decay rate $\alpha\in[0,1]$:
-$b_{G,p}\leftarrow\alpha b_{G,p}+(1-\alpha)R_{p}(\tau)$ (update baseline) (8)
-$A_{G,p}(\tau)=R_{p}(\tau)-b_{G,p}$ (compute advantage) (9)
-This provides better variance reduction than a global baseline by accounting for role-specific asymmetries, such as the first-move advantage. The variance-reduced policy gradient becomes:
+RAE maintains separate baselines $b_{G,p}$ for each game $G\in\mathcal{G}$ and role $p\in\{0,1\}$, estimating expected return $\mathbb{E}[R_{p}(\tau)]$ for that role in that game.
+We update these baselines using exponential moving average with decay rate $\alpha\in[0,1]$:
+$b_{G,p}$ $\leftarrow\alpha b_{G,p}+(1-\alpha)R_{p}(\tau),\quad A_{G,p}(\tau)=R_{p}(\tau)-b_{G,p}$ (2)
+This provides better variance reduction than a global baseline by accounting for role-specific asymmetries. The variance-reduced policy gradient becomes:
 $\nabla_{\theta}J_{\text{SPIRAL}}(\theta)=\mathbb{E}_{G\sim\mathcal{G}}\mathbb{E}_{\tau\sim\pi_{\theta}\times\pi_{\theta}|G}\left[\sum_{p\in\{0,1\}}\sum_{t\in T_{p}}A_{G,p}(\tau)\cdot\nabla_{\theta}\log\pi_{\theta}(y_{t}^{(p)}|s_{t},p,G)\right]$ (3)
-By centering returns around role-specific expectations, RAE ensures that gradient updates reflect genuine learning signal rather than inherent advantages of certain player positions.
-Note that we do not normalize by response length to avoid length bias (Liu et al., 2025b).
+By centering returns around role-specific expectations, RAE ensures gradient updates reflect genuine learning signal rather than inherent positional advantages.
+We do not normalize by response length to avoid length bias (Liu et al., 2025c).
 The complete procedure is in Algorithm 1.
-4.4 Implementation
+Implementation.
 To implement SPIRAL, we develop a truly online multi-agent, multi-turn RL system for finetuning LLMs.
@@
 We instantiate actors to execute the self-play loop, using vLLM (Kwon et al., 2023) for efficient model inference and TextArena (Guertler et al., 2025) to simulate the language games.
-The resulting multi-turn, multi-game self-play experiences are used to update the LLM via policy gradient methods (Sutton and Barto, 2018), incorporating our proposed Role-conditioned Advantage Estimation in the collocated learner.
-5 Experiment Setup
-We design experiments to carefully test how training on different zero-sum games can improve reasoning capabilities that generalize to novel downstream math and reasoning tasks, as well as novel zero-sum games.
-This section introduces the selection of game environments (Sec.
-5.1) and the training (Sec.
-5.2) and evaluation (Sec.
-5.3) settings.
-We investigate four key questions:
-  1. 1.
-    RQ1: Can self-play on zero-sum games improve math and general reasoning capabilities?
-We test whether competitive environment alone can develop transferable reasoning capabilities without domain-specific training data, and analyze emergent patterns to understand transfer mechanisms.
-  2. 2.
-    RQ2: Does self-play’s automatic curriculum outperform fixed-opponent training?
-We evaluate whether the automatic curriculum with adaptive difficulty generated through self-play produces more robust reasoning than training against static opponents.
-  3. 3.
-    RQ3: Do different games develop specialized reasoning skills?
-We test whether distinct game mechanics cultivate complementary cognitive abilities and whether multi-game training creates synergistic benefits.
-  4. 4.
-    RQ4: Is Role-Conditioned Advantage Estimation (RAE) essential for stable self-play training?
-We ablate our variance reduction technique to determine whether RAE is necessary to stabilize training and prevent thinking collapse in zero-sum self-play with shared parameters.
-5.1 Game Environments
-We select three games from TextArena (Guertler et al., 2025) with distinct cognitive requirements:
-TicTacToe (Spatial Reasoning) (Beck, 2008).
-A 3×3 grid game requiring pattern recognition and adversarial planning.
-Players must identify winning configurations, block opponent threats, and plan multi-step forcing sequences.
-We hypothesize these skills transfer to geometric problem-solving and spatial visualization tasks.
-Tic-Tac-Toe is a deterministic, perfect-information game, which isolates pure strategic reasoning from uncertainty management.
-Kuhn Poker (Probabilistic Reasoning) (Kuhn, 1950).
-A minimal poker variant with three cards (Jack, Queen, King) where players bet with hidden information.
-Success requires probability calculation, opponent modeling, and decision-making under uncertainty.
-We expect these capabilities to transfer to problems involving probability, expected value, and strategic uncertainty.
-Simple Negotiation (Strategic Optimization) (Nash, 1950).
-A resource trading game where two players exchange Wood and Gold with opposing valuations to maximize portfolio value.
-Success requires multi-step planning, theory of mind to model opponent preferences, and strategic communication through proposals and counteroffers.
-We hypothesize these skills transfer to optimization problems, resource allocation tasks, and multi-constraint reasoning that requires balancing competing objectives.
-Please see Fig.
-7 in App.
-B for example environment observations.
-5.2 Training Settings
-We employ Qwen3-4B-Base (Yang et al., 2025) as the base model for all experiments.
-The training spans 400 steps with 128 samples per step, totaling 51,200 game transitions.
-We use a sampling temperature $\tau=1.0$ to collect game experiences and Adam (Kingma and Ba, 2014) with a constant learning rate of $1\times 10^{-6}$ to optimize the model.
-The training batch size is set as 128 and the discount factor $\gamma=1.0$ (no discounting).
-We use $\alpha=0.95$ for the exponential moving average decay of role-specific baselines.
-Detailed hyperparameters are listed in Table 7 in App.
-B.
-Each experiment takes around 25 hours for full parameter tuning on 8 H100 GPUs with our distributed actor-learner architecture.
-5.3 Evaluation Metrics
-We track three key metrics throughout the training process: the frequency of invalid moves, the game lengths and the win rates against Gemini-2.0-Flash-Lite [^3] Accessed via https://openrouter.ai/google/gemini-2.0-flash-lite-001., to monitor the learning dynamics and determine the best models.
-To evaluate the trained models, we test their win rates against a fixed opponent on both training and unseen games and standard reasoning benchmarks, as detailed below.
-Out-of-Distribution Generalization to Unseen Games.
-We evaluate transfer learning on seven unseen games that test whether learned skills generalize beyond training domains: Snake and Connect Four (spatial reasoning), Pig Dice and Liar’s Dice (probabilistic reasoning), Truth and Deception (strategic optimization).
-These games specifically probe whether spatial reasoning from TicTacToe, probabilistic reasoning from Kuhn Poker, and strategic optimization from Simple Negotiation transfer to novel game mechanics.
-We report the average win rate of our models against Gemini-2.0-Flash [^4] Accessed via https://openrouter.ai/google/gemini-2.0-flash-001.
-across 512 independent games, with the starting player randomized in each match to eliminate any first-move advantage.
-Transfer to Standard Reasoning Benchmarks.
-To investigate whether the reasoning abilities developed through gameplay could transfer to non-game contexts, we evaluate our models on a suite of established benchmarks.
-For mathematical reasoning, we use MATH500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME24, AIME25, and AMC23 datasets, which cover a wide range of topics including algebra, geometry, and competitive mathematics.
-For general reasoning, we utilize GPQA (Rein et al., 2024), which consists of graduate-level science questions, and MMLU-Pro (Wang et al., 2024), a benchmark for multidisciplinary knowledge.
-All evaluations on these benchmarks are conducted in a zero-shot setting [^5] Except for the base model, for which we provide few-shot examples that follow the Qwen3 Report settings.
-to determine if game-induced reasoning could be successfully transferred to general problem-solving.
-We use sampling temperature $0.6$, top-p $0.95$ and $4$ rollouts for all evaluations.
-Reasoning Pattern Analysis.
-We analyze the cognitive patterns that emerge during training and their transfer to mathematical reasoning.
-Using GPT-4.1 (gpt-4.1-2025-04-14(OpenAI, 2025) as LLM-as-Judge (Zheng et al., 2023), we classify reasoning traces from 290 game trajectories and 46,792 math problem solutions into three core patterns: Case-by-Case Analysis (systematic enumeration of scenarios), Expected Value Calculation (probabilistic decision-making), and Pattern Recognition (identifying regularities and structures).
-The complete prompts used for pattern discovery, classification, and transfer analysis are provided in App.
-D.
-We track the evolution of these patterns across training checkpoints (early: step 0, mid: step 128, late: step 400) in both game and math domains to measure transfer rates.
-This analysis reveals which game-induced reasoning strategies generalize to academic problem-solving and explains the mechanism behind cross-domain transfer.
-6 Experimental Results and Findings
-6.1 RQ1: Can self-play on zero-sum games improve math and general reasoning capabilities?
-Table 1 presents our main results.
-SPIRAL, trained exclusively on Kuhn Poker through self-play, achieves notable improvements on both mathematical and general reasoning benchmarks.
-Notably, it even surpasses a model that was fine-tuned on 25,000 winning game trajectories from Qwen3-32B playing against itself.
-Table 1: Reasoning benchmark performance. SPIRAL improves reasoning without any domain-specific training data.
-Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average
-Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 32.9
-SFT (Qwen3-32B Distill) 76.0 16.7 13.3 38.4 50.0 31.2 33.0 48.8 39.7
-SPIRAL (Ours) 76.4 13.3 10.0 38.4 57.5 42.4 37.0 57.7 41.6
-Improvement +10.6 +3.3 +6.7 +5.1 +7.5 +18.1 +6.4 +10.5 +8.7
+The resulting multi-turn, multi-game self-play experiences are used to update the LLM via policy gradient methods (Sutton & Barto, 2018), incorporating our proposed Role-conditioned Advantage Estimation in the collocated learner.
+4 Experimental Results
+Table 1: Reasoning benchmark performance.
+The “-Kuhn” suffix denotes fine-tuning solely on a single game (Kuhn Poker), while the “-Multi” suffix indicates fine-tuning on all three games.
+SPIRAL improves reasoning without any domain-specific training data.
 Few-shot evaluation following Qwen3 technical report.
-The results show consistent improvements across mathematical reasoning (+10.6% MATH500, +6.7% AIME’25, +7.5% AMC-23), and general reasoning (+6.4% GPQA, +10.5% MMLU-Pro).
-Notably, SPIRAL achieves these gains despite never seeing a single mathematical equation, academic problem, or domain-specific training example during its entire training process.
-The 18.1% improvement on Minerva Math and overall 8.7% gain across math benchmarks demonstrate that competitive game environments successfully force discovery of reasoning patterns that transfer broadly across domains.
-6.1.1 Case Study: How Game Strategies Transfer to Mathematical Reasoning
-To understand this surprising transfer, we use GPT-4.1 (gpt-4.1-2025-04-14) to analyze 290 game trajectories and 46,792 math solutions across training checkpoints, identifying three core reasoning patterns that emerge during gameplay and transfer to math reasoning domains.
-Pattern Discovery and Examples. We first illustrate what these patterns look like in both game and mathematical contexts:
-Table 2: Representative reasoning patterns and their manifestation in games and mathematics. We use an LLM-as-a-judge framework to automatically evaluate the frequency of these patterns over training.
+Model Math500 AIME24 AIME25 Olympiad AMC-23 Minerva GPQA-D MMLU-Pro Average
+Qwen3-4B-Base $73.4$ $9.6$ $6.2$ $33.3$ $42.4$ $29.4$ $30.6$ $47.2$ $34.0$
+   + SFT-Kuhn $74.0$ $11.0$ $10.4$ $36.7$ $48.6$ $36.8$ $33.0$ $48.8$ $37.4$
+   + SFT-Multi $74.2$ $13.7$ $11.7$ $37.6$ $51.1$ $40.1$ $37.8$ $51.3$ $39.7$
+   + SPIRAL-Kuhn (Ours) $76.4$ $18.2$ $15.6$ $38.4$ $61.2$ $42.4$ $37.0$ $57.7$ $43.4$
+   + SPIRAL-Multi (Ours) 78.278.2+4.8 19.719.7+10.1 13.313.3+7.1 41.841.8+8.5 61.661.6+19.2 42.642.6+13.2 40.140.1+9.5 58.558.5+11.3 44.544.5+10.5
+Qwen3-8B-Base $77.0$ $12.1$ $11.2$ $33.5$ $50.6$ $38.2$ $38.0$ $55.7$ $39.5$
+   + SFT-Multi $82.8$ $19.9$ $15.6$ $45.9$ $63.5$ $40.8$ $41.6$ $58.8$ $46.1$
+   + SPIRAL-Multi (Ours) 86.686.6+9.6 26.226.2+14.1 16.816.8+5.6 49.649.6+16.1 65.265.2+14.6 46.346.3+8.1 44.644.6+6.6 61.161.1+5.4 49.649.6+10.1
+Octothinker-8B-Base $65.6$ $1.7$ $0.5$ $26.6$ $33.5$ $25.7$ $22.1$ $30.8$ $25.8$
+   + SFT-Multi $66.0$ $3.3$ $3.8$ $23.9$ $31.0$ $23.8$ $24.9$ $39.1$ $27.0$
+   + SPIRAL-Multi (Ours) 68.668.6+3.0 5.35.3+3.6 4.84.8+4.3 33.733.7+7.1 43.243.2+9.7 32.032.0+6.3 33.833.8+11.7 49.349.3+18.5 33.833.8+8.0
+Llama-3.1-8B-Instruct $46.4$ $4.6$ $0.7$ $13.8$ $23.3$ $22.8$ $30.2$ $49.1$ $23.9$
+   + SFT-Multi $51.8$ $4.6$ $0.7$ 19.1 $23.3$ $21.7$ $30.0$ $48.9$ $25.0$
+   + SPIRAL-Multi (Ours) 49.849.8+3.4 4.94.9+0.3 1.81.8+1.1 17.317.3+3.5 26.026.0+2.7 24.624.6+1.8 32.232.2+2.0 50.450.4+1.3 25.925.9+2.0
+DeepSeek-Distill-Qwen-7B $90.8$ $53.0$ $39.5$ $56.9$ $89.3$ $48.2$ $48.6$ $57.1$ $60.4$
+   + SFT-Multi $91.8$ $49.3$ $36.6$ $52.4$ $88.2$ $48.2$ $44.5$ $55.6$ $58.3$
+   + SPIRAL-Multi (Ours) 93.093.0+2.2 54.154.1+1.1 40.840.8+1.3 57.957.9+1.0 89.389.3+0.0 51.151.1+2.9 49.649.6+1.0 58.958.9+1.8 61.861.8+1.4
+We evaluate SPIRAL across diverse model architectures and game environments to understand how self-play develops transferable reasoning capabilities.
+We train on three games from TextArena (Guertler et al., 2025): TicTacToe (spatial reasoning), Kuhn Poker (probabilistic reasoning), and Simple Negotiation (strategic optimization).
+Models include Qwen3-4B/8B-Base (Yang et al., 2025), Llama-3.1-8B-Instruct (Dubey et al., 2024), and Octothinker-8B-Base (Wang et al., 2025a).
+Training spans 400 steps with 128 samples per step on 8 H100 GPUs, using Adam optimizer with learning rate $1\times 10^{-6}$ and temperature 1.0.
+We evaluate on eight reasoning benchmarks (MATH500, OlympiadBench, Minerva Math, AIME24/25, AMC23, GPQA-Diamond, MMLU-Pro) and seven out-of-distribution games.
+Complete implementation details are in Appendix D.
+Self-play on games transfers to improve reasoning.
+The central results of this paper are shown in Table 1, which demonstrates that multi-game SPIRAL training achieves up to 10.5% improvement on reasoning benchmarks without domain-specific data.
+with Qwen3-4B-Base improving from 34.0% to 44.5% average performance (+10.5+10.5%).
+We also compare with supervised fine-tuning (SFT) on 25,000 expert game trajectories, generated by Qwen3-32B models, which improves performance on several benchmarks including AIME24 and AIME25, revealing that games themselves contain skills relevant to reasoning.
+However, SPIRAL consistently outperforms SFT across all 8 benchmarks, demonstrating that self-play discovers more effective reasoning strategies than imitating expert demonstrations.
+4.1 Understanding Why SPIRAL Works
+Figure 4: Evolution of reasoning patterns during SPIRAL training and their transfer to mathematical reasoning.
+We track three core reasoning patterns (Pattern Recognition, Expected Value Calculation, and Case-by-Case Analysis) across 290 game trajectories and 46,792 math solutions.
+Left: Patterns increase via training on games, with Expected Value Calculation reaching 78% by late training.
+Middle: These patterns transfer to mathematical reasoning with varying effectiveness: Case-by-Case Analysis maintains high transfer (72% to 71%), Pattern Recognition shows amplification (35% to 45%), while Expected Value Calculation transfers more selectively (78% to 28%).
+Right: Math benchmark scores improve from 31.2 to 39.6 as these reasoning patterns develop, demonstrating that game-learned strategies enhance mathematical problem-solving capabilities.
+Emergent reasoning patterns learned on games transfer to math questions.
+To understand how games can improve reasoning performance, we analyzed chain-of-thought patterns using GPT-4.1 (OpenAI, 2025) to classify 290 game trajectories and 46,792 math solutions.
+Table 2 illustrates three core patterns that emerge during gameplay and transfer to mathematics.
+Figure 4 tracks their evolution.
+In the late training stage, the transfer from games to math is clear: Case-by-Case Analysis transfers near-perfectly (72% to 71%), Pattern Recognition amplifies (35% to 45%), while Expected Value Calculation transfers selectively (78% to 28%).
+The success of this transfer may stem from two factors.
+This transfer likely works because evolving opponents prevent memorization while games teach pure reasoning strategies without requiring mathematical knowledge.
+Table 2: Representative reasoning patterns and their manifestations in games and mathematics. We use an LLM-as-a-judge framework to evaluate the frequency of these patterns over training.
 Pattern Game Context (Kuhn Poker) Math Transfer
-Case-by-Case Analysis <think>
+Case-by-Case Analysis <output>
 Case 1 - Fold: You lose 1 chip. You have 2 chips now.
 Case 2 - Call: You have 0% chance of winning, so you will have 1 chip if you lose.
-Since losing 1 chip (fold) is better than losing 2 chips (call and lose), the best action is to fold.
-</think>
-<think> To solve $m^{2}+2\cdot 3^{n}=m(2^{n+1}-1)$:
+Since losing 1 chip (fold) is better than losing 2 chips (call and lose), the best action is to $\boxed{fold}$.
+</output>
+<output> To solve $m^{2}+2\cdot 3^{n}=m(2^{n+1}-1)$:
 Case $n=0$: $m^{2}+2=m$, no real solutions
@@
 Case $n=3$: $m^{2}+54=15m$ gives $m=6,9$ (valid)
-Case $n\geqslant 4$: No solutions (exponential dominance) </think>
-Expected Value Calculation <think> With card J against opponent’s bet:
+Case $n=4$: $m^{2}+162=31m$, no integer solutions
+Case $n=5$: $m^{2}+486=63m$ gives $m=\boxed{9,54}$ (valid) </output>
+Expected Value Calculation <output> With card J against opponent’s bet:
 EV(call) = P(win) $\times$ 2 - P(lose) $\times$ 2
@@
 EV(fold) = -1 (lose ante)
-Since EV(fold) > EV(call), fold is optimal. </think>
-<think> For average of $A+BC+DEF$ over permutations:
+Since EV(fold) >> EV(call), $\boxed{fold}$ is optimal. </output>
+<output> For average of $A+BC+DEF$ over permutations:
 E[A] = $\frac{\sum a_{i}}{6}=\frac{-2}{6}=-\frac{1}{3}$
@@
 E[DEF] = $(E[D])(E[E])(E[F])=(-\frac{1}{3})^{3}=-\frac{1}{27}$
-Total = $-\frac{1}{3}+\frac{1}{9}-\frac{1}{27}=-\frac{7}{27}$ </think>
-Pattern Recognition <think> Previous rounds: Player 0 had K both times and won.
+Total = $\frac{1}{3}+\frac{1}{9}-\frac{1}{27}=\boxed{-\frac{7}{27}}$ </output>
+Pattern Recognition <output> Previous rounds: Player 0 had K both times and won.
 Pattern identified: Player 0 likely has strong cards or bluffs consistently.
 Given I have J (weak), betting might exploit their aggressive calling pattern.
-</think> <think> Sum of three consecutive integers $(n-1)+n+(n+1)=3n$.
+</output> <output> Sum of three consecutive integers (n1)+n+(n+1)=3n(n-1)+n+(n+1)=3n.
 Pattern recognized: sum is always divisible by 3.
 For perfect cube: $3n=k^{3}$, so $k$ must be divisible by 3.
-Smallest: $k=3\Rightarrow 3n=27$ </think>
-Pattern Evolution and Transfer Analysis. Using GPT-4.1 to classify reasoning strategies, we tracked how these patterns develop during training and transfer to mathematics:
-Figure 4: Evolution of reasoning patterns during SPIRAL training and their transfer to mathematical reasoning.
-We track three core reasoning patterns (Pattern Recognition, Expected Value Calculation, and Case-by-Case Analysis) across 290 game trajectories and 46,792 math solutions.
-Left: In the game domain, all patterns show substantial growth, with Expected Value Calculation reaching 78% by late training.
-Middle: These patterns transfer to mathematical reasoning with varying effectiveness: Case-by-Case Analysis maintains high transfer (72% to 71%), Pattern Recognition shows amplification (35% to 45%), while Expected Value Calculation transfers more selectively (78% to 28%).
-Right: Math benchmark scores improve from 31.2 to 39.6 as these reasoning patterns develop, demonstrating that game-learned strategies enhance mathematical problem-solving capabilities.
-Fig. 4 shows how reasoning patterns evolve in both game and math domains across training checkpoints. These patterns develop at different rates and transfer with varying effectiveness:
-Case-by-Case Analysis: We find that the meta-cognitive skill of systematic enumeration, or breaking problems down into cases, transfers highly effectively from strategic games to math problems (as shown in Fig.
-4).
-We hypothesize this is because it represents a domain-agnostic way of structuring thinking that generally improves reasoning performance.
-Whether analyzing opponent possibilities in Poker or solution branches in mathematics, the core skill is the same.
-Expected Value Calculation: We find that training on games increases the use of expected value calculation from 15% to 78% in the games themselves, and from 12% to 28% in the math domain.
-While game-specific probabilistic reasoning is more frequent in games than in the math benchmark (likely because most math problems lack explicit decision-theoretic structure), we find that this form of reasoning benefits probability and optimization problems where EV concepts directly apply.
-Pattern Recognition: Interestingly, this form of reasoning shows an amplification effect, where the proportion evident in the math domain actually exceeds that of the game domain.
-We hypothesize that because mathematics inherently requires pattern recognition, game training enhances an already-present mathematical skill which is then deployed even more frequently when the model is prompted to answer a math problem.
-The combined training produces stronger pattern recognition (45%) than games alone require (35%).
-Transfer Mechanism. We believe three factors enable this cross-domain transfer:
-1. Competitive pressure strips away memorization: Self-play opponents continuously evolve, forcing models to develop genuine reasoning rather than pattern matching.
-2. Games isolate pure reasoning: Without domain complexity, games teach fundamental cognitive operations (enumeration, evaluation, synthesis) that generalize effectively.
-3. Structured output bridges domains: The <think> format learned in games provides a reasoning scaffold that models reuse for mathematical problems.
-Finding 1: Zero-sum games teach transferable reasoning without domain data Self-play on Kuhn Poker improves math and general reasoning benchmarks despite never seeing benchmark related problems.
-This works because gameplay forces discovery of three reasoning patterns: Case-by-Case Analysis, Expected Value Calculation, and Pattern Recognition.
-These skills are developed during competing against an evolving opponent.
-6.2 RQ2: Does self-play’s automatic curriculum outperform fixed-opponent training?
-Although self-play promises potentially infinite curriculum generation, it remains unclear whether the evolving challenge actually produces better learning than simply training against fixed opponents.
-We test whether adaptive difficulty through self-play leads to more robust reasoning development.
-Experimental Setup.
-We compare four training settings on the Kuhn Poker environment: (1) Self-Play: the model trains against continuously updating copies of itself; (2) Random Opponent: the opponent always provides valid actions but poses minimal strategic challenge; (3) Mistral Opponent: using Mistral-Small-3 (Mistral, 2025) as a fixed intermediate-level opponent; and (4) Gemini Opponent,: using Gemini-2.0-Flash-Lite (Gemini Team, Google, 2025) as a fixed strong opponent.
-All variants use identical hyperparameters, infrastructure, and 400 training steps.
-Results Analysis.
-Figure 5 reveals two distinct failure modes in fixed-opponent training:
-Figure 5: Performance comparison of self-play training and fixed-opponent baselines.
-All evaluations are averaged over multiple games/benchmarks (see Sec.
-5.3).
-Mistral Opponent refers to against Mistral-Small-3; Gemini Opponent refers to against Gemini-2.0-Flash-Lite.
-1.
-Curse of Turns in Format Learning: Training against random opponents leads to complete collapse.
-Although these rule-based random agents offer no strategic challenge, they always produce valid moves.
-Consequently, the model must generate correctly formatted, valid actions at every turn throughout the trajectory in order to receive any positive reward.
-However, the probability of producing a fully valid trajectory decreases exponentially with the episode length, making it extremely difficult to explore and learn the correct action format over long horizons.
-2.
-Exploitation of Static Strategies: Fixed model-based opponents (Mistral, Gemini) facilitate format learning more effectively than random opponents, as they occasionally produce invalid moves that allow the agent to win and receive reward signals.
-However, we find that training against these static opponents often leads to overfitting to their fixed strategies, resulting in poor generalization to math and general reasoning benchmarks (middle and right of Fig.
-5).
-Continuous Adaptation Through Self-Play.
-Self-play uniquely maintains learning momentum throughout training and generalizes better because self-play automatically adjusts difficulty.
-As shown in Table 3, training against a fixed opponent (Gemini) initially gets zero win rate (thus no positive reward), leading to no effective learning before step 128 (see left of Fig.
-5); at step 384 the agent learns to outperform the fixed opponent with $62.5\%$ win rate, gradually exploiting its static strategies.
-On the other hand, against previous versions of itself, self-play maintains 50-52% win rates throughout training, confirming the opponent evolves to match current capabilities.
+Smallest: $k=\boxed{3}\Rightarrow 3n=27$ </output>
+Adaptive curriculum beats static opponents.
+Self-play creates an automatic curriculum that adapts to model capabilities.
+Figure 5 compares self-play against fixed opponents (Random, Mistral-Small-3, Gemini-2.0-Flash-Lite [^1] Accessed via https://openrouter.ai/google/gemini-2.0-flash-lite-001.).
 Table 3: Win rates at different training stages of Gemini Opponent and Self-Play vs its opponent.
@@
 Step 384 62.5% 50.9%
-The reasoning transfer results are even more striking.
-Self-play achieves 40% on math reasoning and 45% on general reasoning, outperforming the best fixed opponent (Gemini) by 5 and 3 percentage points respectively.
-The relative improvement demonstrates that the diverse strategies required by an evolving opponent create more generalizable reasoning patterns than exploiting static weaknesses.
-Finding 2: Self-play creates adaptive curriculum that outperforms fixed opponents Self-play achieves better reasoning transfer than any fixed-opponent training.
-Rule-based random opponents cause complete collapse due to the curse of turns in format learning.
-Training against model-based opponents plateaus once the agent finds winning counter-strategies.
-Self-play avoids both failures by continuously adjusting difficulty: the opponent improves as the agent improves, forcing ongoing adaptation rather than exploitation of static weaknesses.
-6.3 RQ3: Do different games develop specialized reasoning skills?
-We investigate whether distinct game mechanics cultivate different cognitive abilities.
-If games truly serve as reasoning gymnasiums, then training on TicTacToe should develop different skills than training on Kuhn Poker or Simple Negotiation.
-Furthermore, we test whether combining multiple games produces synergistic benefits beyond the sum of individual training.
-Experimental Design.
-To isolate game-specific skill development, we train separate models exclusively on each game using identical hyperparameters and training duration.
-We then evaluate these specialists in three ways: (1) performance on their training game, (2) transfer to out-of-distribution games, and (3) generalization to math and general reasoning benchmarks.
-This design reveals whether games develop narrow tactics or transferable cognitive capabilities.
-For game performance evaluation, we use two complementary approaches: head-to-head competition between specialists to directly compare learned strategies, and performance against a fixed strong opponent (Gemini-2.0-Flash-Lite) to measure absolute skill levels.
-This dual evaluation reveals both relative strengths between specialists and their absolute capabilities.
-Head-to-Head Competition Reveals Specialized Strategies.
-We first examine how specialists perform when competing directly against each other.
-This head-to-head evaluation uses games that isolate specific cognitive skills: we pair each training game with an unseen game requiring similar abilities (Snake for spatial reasoning like TicTacToe, Pig Dice for probabilistic reasoning like Kuhn Poker, and Truth and Deception for strategic optimization like Simple Negotiation).
+Random opponents cause collapse: although they provide randomized rewards with positive expected value, similar to spurious rewards that might upweight certain base model behaviors (Shao et al., 2025), we observe this is insufficient to improve performance.
+Fixed model opponents enable initial learning but plateau once exploitable strategies are found.
+That fixed opponents like Gemini yield smaller gains reveals the effects are not merely from learning game mechanics or spurious rewards (Shao et al., 2025), but specifically from the adaptive curriculum.
+Unlike static baselines, self-play’s continuously evolving challenge forces genuine reasoning development rather than exploitation of static patterns.
+Table 3 confirms this: self-play maintains 50-52% win rates while fixed-opponent training rises from 0% to 62.5%, indicating exploitation rather than continued learning.
+Figure 5: Performance comparison of self-play training and fixed-opponent baselines.
+All evaluations are averaged over multiple games/benchmarks (see Sec.˜D.3).
+Mistral Opponent refers to against Mistral-Small-3; Gemini Opponent refers to against Gemini-2.0-Flash-Lite.
+Different games develop complementary skills.
+Each game cultivates distinct cognitive abilities that transfer to related domains; specifically, we selected TicTacToe to target spatial reasoning, Kuhn Poker for probabilistic inference, and Simple Negotiation for strategic optimization.
+Table 4 tests how well agents trained on a specific game (‘specialists’) transfer to novel out-of-distribution (OOD) games.
+We find specialists transfer effectively to similar out-of-distribution games: TicTacToe specialists achieve 56.0% on Snake (spatial), Poker specialists dominate Pig Dice at 91.7% (probabilistic), Negotiation specialists win 55.8% on Truth and Deception (strategic).
+Multi-game training combines these skills synergistically, as shown in Table 5, which shows win-rate against Gemini-2.0-Flash [^2] Accessed via https://openrouter.ai/google/gemini-2.0-flash-001..
+Multi-game agents achieve 59.5% average performance, outperforming all single-game specialists (best: 52.9%), using a fixed budget of 400 training steps determined by convergence analysis.
 Table 4: Game specialists excel at both their training games and unseen games requiring similar cognitive skills.
@@
 Negotiation Specialist 40.5% 40.2% 62.7% 41.0% 1.1% 55.8%
-Models trained exclusively on TicTacToe excel at spatial pattern recognition, achieving 57.5% win rate against other specialists on their training game.
-When tested on Snake, an unseen spatial reasoning game, they maintain 56.0% performance, demonstrating robust transfer.
-Kuhn Poker specialists develop probabilistic reasoning, dominating their training game (64.2%) and remarkably achieving 91.7% on Pig Dice.
-Simple Negotiation training produces models skilled in strategic optimization (62.7% on training game, 55.8% on Truth and Deception).
-Multi-Game Training Creates Synergistic Benefits.
-While head-to-head competition reveals specialized strategies, we also evaluate absolute performance against a strong fixed opponent (Gemini-2.0-Flash-Lite) across both training games and randomly sampled out-of-distribution games.
-The multi-game model demonstrates broader capabilities that often exceed specialist performance:
 Table 5: Multi-game training achieves competitive performance across all training games while excelling at novel composite challenges.
-All win rates shown are against Gemini-2.0-Flash-Lite as a fixed opponent.
+All win rates shown are against Gemini-2.0-Flash as a fixed opponent.
 The multi-game model outperforms all specialists on average, demonstrating that diverse game training develops more flexible reasoning.
 Training Games OOD Games
-Model TicTacToe KuhnPoker Simple Negotiation Liar’s Dice Connect Four Snake Average
-Base Model 17.5% 21.5% 15.6% 16.7% 20.5% 20.6% 18.7%
+Model TicTacToe KuhnPoker Simple Negotiation Snake Pig Dice Truth and Deception Average
+Base Model 17.5% 21.5% 15.6% 7.8% 0.2% 49.6% 18.7%
+Random Policy 24.5% 31.3% 8.2% N/A N/A N/A N/A
+Instruct Model 52.7% 48.5% 46.2% 25.2% 97.6% 75.5% 57.6%
 Single-Game Specialists
-      TicTacToe Specialist 56.6% 24.4% 30.5% 31.4% 30.8% 30.7% 34.1%
-      Kuhn Poker Specialist 31.0% 48.5% 28.7% 24.9% 39.1% 33.9% 34.4%
-      Simple Negotiation Specialist 27.7% 16.8% 39.1% 12.3% 32.0% 34.8% 27.1%
-Multi-Game Model 54.3% 53.9% 33.2% 51.4% 39.3% 37.2% 44.9%
-The power of multi-game training becomes evident in performance on novel games.
-On Liar’s Dice, individual specialists struggle, achieving only 24.9% (Kuhn Poker) and 12.3% (Simple Negotiation) against Gemini-2.0-Flash-Lite.
-However, the multi-game model achieves 51.4%, demonstrating emergent capabilities beyond any single training game.
-This pattern of multi-game superiority holds across the out-of-distribution games, with the multi-game model achieving the highest average performance (44.9%) compared to all specialists.
-Multi-Game Training Enhances Even Strong Reasoning Models.
-To test whether game-based skill development is limited to weaker models, we apply multi-game SPIRAL training to DeepSeek-R1-Distill-Qwen-7B, a model that already achieves 59.7% average on reasoning benchmarks (see Sec.
-C.2 for comprehensive results across all training configurations):
-Table 6: Multi-game SPIRAL training improves reasoning capabilities across model scales. Even strong reasoning models benefit from game-based skill development.
-Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average
-Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 33.1
-+ SPIRAL (Multi-Game) 74.2 20.0 16.7 39.4 55.0 38.6 36.9 57.7 42.3
-DeepSeek-Distill-Qwen-7B 90.8 46.7 36.7 56.9 92.5 48.2 48.6 57.1 59.7
-+ SPIRAL (Multi-Game) 93.0 43.3 46.7 57.9 92.5 51.1 49.6 58.9 61.7
-Multi-game SPIRAL improves Qwen3-4B by 9.2 percentage points (33.1% to 42.3%) and enhances the already-strong DeepSeek-Distill-Qwen-7B model by 2.0 points (59.7% to 61.7%).
-This demonstrates that game-based reasoning skills provide value even for models that already excel at mathematical reasoning, suggesting that games teach complementary cognitive abilities not captured by traditional training.
-Finding 3: Different games develop specialized skills that work together Each game builds distinct abilities that transfer to similar out-of-distribution games.
-Multi-game training combines these specialized skills, achieving superior performance on novel games that individual specialists struggle with.
-This synergistic training improves reasoning benchmarks even for already-strong models, demonstrating that games teach complementary cognitive skills.
-6.4 RQ4: Is Role-Conditioned Advantage Estimation (RAE) essential for stable self-play training?
-We ablate our variance reduction technique to determine whether RAE is necessary to stabilize training and prevent thinking collapse in zero-sum self-play with shared parameters.
-Figure 6 reveals that without proper variance reduction, self-play training catastrophically fails:
-1.
-Thinking Collapse: The most striking failure occurs in response length (bottom right).
-Without RAE, reasoning traces plummet from 3,500 to near-zero characters after 200 steps.
-Models quite literally stop thinking, generating minimal outputs like <think></think><answer>
-boxed{bet}</answer>
-. This collapse coincides with gradient norms approaching zero, indicating the model has converged to a degenerate policy.
-2.
-Performance Degradation: Without RAE, the model learns solely to play the game while abandoning CoTs, leading to generalization failure.
-As a result, the math reasoning performance crashes from 35% to 12% at around step 150 (a 66% relative decrease), and the general reasoning drops from 44% to 40%.
-3.
-Training Instability: Policy gradient norms without RAE exhibit higher initial values with spikes, then collapses to near-zero after step 200.
-In contrast, RAE maintains stable gradient norms of around 0.1 throughout training, enabling continuous learning.
-This variance reduction is crucial for maintaining stable reasoning.
-High-variance gradients cause models to oscillate between extreme strategies, eventually converging to degenerate short responses that minimize variance but abandon reasoning entirely.
+     TicTacToe Specialist 56.6% 24.4% 30.5% 28.1% 97.6% 79.9% 52.9%
+     Kuhn Poker Specialist 31.0% 48.5% 28.7% 27.7% 98.8% 81.6% 52.7%
+     Simple Negotiation Specialist 27.7% 16.8% 39.1% 26.4% 98.6% 82.8% 48.6%
+Multi-Game Model 54.3% 53.9% 33.2% 31.6% 99.8% 84.0% 59.5%
 Figure 6: Training dynamics comparing REINFORCE with RAE (orange) versus vanilla REINFORCE (gray).
 RAE maintains stable performance across all metrics while vanilla REINFORCE suffers catastrophic thinking collapse.
-Top row: Performance on game playing, math reasoning, and general reasoning benchmarks.
-Bottom row: Policy gradient norm shows training instability without RAE; response length reveals thinking collapse where models stop generating reasoning traces.
-Finding 4: RAE prevents models from abandoning reasoning during self-play Without RAE, models began to truncate their reasoning processes after 200 steps, generating empty reasoning traces like <think></think>.
-This thinking collapse causes math performance to crash while gradients become unstable then drop to zero.
-RAE maintains stable gradients throughout training, allowing models to keep generating substantive reasoning.
-Self-play alone is not enough; proper variance reduction is essential for maintaining reasoning in competitive games.
+Left: Response length reveals thinking collapse where models stop generating reasoning traces; Right: Performance on general reasoning benchmarks.
+[…]
 5 Conclusion
-Summary of Contributions.
 We introduced SPIRAL, enabling language models to develop reasoning capabilities through competitive self-play without human-curated data.
 Our technical contributions include a fully online multi-turn MARL system for LLMs and Role-conditioned Advantage Estimation (RAE), which prevents thinking collapse in zero-sum games.
-Our empirical results show that training on Kuhn Poker alone improves mathematical reasoning by 8.7% average and Minerva Math by 18.1%, surpassing models trained on 25,000 expert demonstrations.
-Different games develop distinct transferable skills, and multi-game training achieves superior performance on both reasoning benchmarks and novel out-of-distribution games, demonstrating synergistic benefits from diverse game training.
-Limitations.
-While eliminating human-curated problems, SPIRAL still requires designed game environments.
-Our experiments use simple games (TicTacToe, Kuhn Poker, Simple Negotiation); scaling to complex environments remains unexplored.
-The computational requirements are substantial (8 H100 GPUs for 25 hours per experiment).
-Performance plateaus after extended training, and our evaluation focuses on academic benchmarks rather than real-world reasoning tasks requiring common sense or ethical judgment.
-Future Directions.
-This work opens several avenues: expanding to cooperative games, incorporating partial observability, and designing games targeting specific reasoning weaknesses.
-Understanding why certain games develop particular skills could enable principled environment design.
-We envision ecosystems of self-improving agents generating increasingly sophisticated challenges, creating autonomous reasoning development beyond human supervision.
-Final Thoughts.
+Empirically, multi-game SPIRAL training improves reasoning benchmarks by up to 10.5% across diverse model architectures, surpassing supervised fine-tuning on 25,000 expert game trajectories.
+Different games develop distinct transferable skills (spatial, probabilistic, strategic) that combine synergistically.
+Analysis reveals that competitive gameplay forces discovery of reasoning patterns (case-by-case analysis, expected value calculation, pattern recognition) that transfer to academic domains.
 SPIRAL demonstrates that simple games can unlock complex reasoning without domain-specific data.
-By harnessing competitive pressure, we create systems that discover their own curricula and continuously improve.
-The transfer from gameplay to mathematics could suggest that intelligence emerges not from sophisticated supervision but could from environmental challenges that force models to think.
-This paradigm shift points toward AI systems that autonomously push reasoning boundaries and continuously evolve through self-play.
-Appendix
-This appendix provides comprehensive details supporting our main findings.
-App.
-A shows how existing SFT and RLVR paradigms adapt to our turn-level MDP formulation.
-App.
-B contains implementation specifics including hyperparameters, infrastructure details, and game environment observations referenced in Section 5.2.
-App.
-C presents extended results tables and detailed pattern evolution analysis that support our case study in Section 6.1.
-App.
-D describes the complete methodology for our reasoning pattern analysis, detailing how we discovered and quantified transfer mechanisms.
-Finally, App.
-E provides comprehensive descriptions of for all game environments used in our experiments, including both training and evaluation games.
-Appendix A Training Paradigms in Turn-Level MDPs
-This section shows how existing training paradigms adapt to our turn-level MDP formulation introduced in Section .
-A.1 Supervised Fine-Tuning (SFT) in Turn-Level MDPs
+Future work could expand to cooperative games, incorporate partial observability, and design games targeting specific reasoning weaknesses.
+Understanding game-skill mappings could enable principled environment design for autonomous reasoning development.
+Acknowledgements
+We thank Xidong Feng and Runji Lin for their helpful discussions and support throughout this project.
+We thank John Schulman for insightful feedback and encouragement.
+We would like to thank Thinking Machine Lab and Modal Lab for providing compute credits that supported our experiments.
+We also thank the TextArena team for maintaining the game environments used in this work.
+This research was supported by the Cooperative AI Foundation, the UW-Amazon Science Gift Hub, Sony Research Award, UW-Tsukuba Amazon NVIDIA Cross Pacific AI Initiative (XPAI), the Microsoft Accelerate Foundation Models Research Program, Character.AI, DoorDash, and the Schmidt AI2050 Fellows program.
+This material is based upon work supported by the Defense Advanced Research Projects Agency and the Air Force Research Laboratory, contract number(s): FA8650-23-C-7316.
+Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of AFRL or DARPA.
+Reproducibility statement
+We provided the experiment code at https://github.com/spiral-rl/spiral.
+We have also provided the training settings in the experiment section (§4) and Appendix D.
+The experiments are run with an 8 H100 GPU cluster, and all calls for proprietary LLMs are via the official API or OpenRouter [^3] https://openrouter.ai/.
+Appendix A Large Language Model Usage
+We used large language models (LLMs) only for language refinement tasks, including grammar checking, phrasing adjustments, and enhancing readability.
+We also used Deep Research (OpenAI, 2025) to assist with related work search.
+Besides these, all scientific ideas, experiments, analyses, and results are the sole contributions of the authors.
+Appendix B Limitations
+Our study, while promising, has several limitations that offer avenues for future research.
+Reliance on Designed Game Environments
+A core limitation is the dependency on engineered game environments.
+Although SPIRAL eliminates the need for human-curated problem datasets, it shifts the dependency to well-designed games.
+The games used in our experiments, such as the Tictactoe and Khun Poker, are relatively simple and feature dense rewards.
+It is an open question how well this approach scales to more complex, open-ended environments with sparse rewards, such as Minecraft or realistic robotics simulations.
+The design of the game environment itself may implicitly encode biases or heuristics that influence the agent’s learned reasoning strategies, potentially limiting their generality.
+Computational Cost and Scalability
+The computational requirements for training are substantial.
+Each experimental run demanded 8 H100 GPUs for approximately 25 hours, which may be prohibitive for many research groups.
+Furthermore, we observed that performance gains began to plateau after extended training periods.
+This suggests that simply scaling up the training duration with the current framework may yield diminishing returns, and more efficient algorithms or architectural improvements are necessary for further progress.
+Evaluation and Transferability
+Our evaluation, while comprehensive, has two key constraints:
+  • Focus on Academic Benchmarks: We primarily assessed reasoning on established academic benchmarks like MATH and GPQA.
+These benchmarks are excellent for measuring formal and scientific reasoning but do not capture the full spectrum of self-play.
+  • Zero-Shot Evaluation: The strict zero-shot evaluation setting tests for direct transfer but may not fully reveal the model’s potential.
+Fine-tuning on a small set of target domain examples could potentially unlock significantly better performance, a possibility not explored in this work.
+Potential for Reward Hacking
+Like many reinforcement learning systems, SPIRAL is susceptible to reward hacking.
+An agent might discover policies that maximize the in-game score without learning the intended underlying reasoning skill.
+For instance, it could exploit a bug in the game physics or find a repetitive, degenerate strategy that succeeds for a narrow set of problems.
+While we did not observe significant instances of this, it remains a risk, especially in more complex environments where robust reward shaping is challenging.
+Appendix C Preliminaries
+This section provides the mathematical foundations and formal definitions underlying the SPIRAL framework, including turn-level MDPs, two-player zero-sum Markov games, and how existing training paradigms adapt to these formulations.
+C.1 Turn-level Markov Decision processes (MDPs).
+Language model training traditionally formulates generation as a token-level MDP (Bellman, 1957; Rafailov et al., 2024) where each action is a single token from vocabulary $\mathcal{V}$.
+For multi-turn reasoning and game-playing, we instead adopt a turn-level MDP formulation $\mathcal{M}=(\mathcal{S},\mathcal{A},T,r,\gamma)$.
+Here, states $\mathcal{S}$ represent complete contexts (e.g., game configurations, problem states, or conversation histories), actions $\mathcal{A}$ are complete responses (containing many tokens), the transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ determines state dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides immediate rewards, and $\gamma\in[0,1]$ is the discount factor.
+The return is defined as the discounted sum of rewards: $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r_{t}$.
+The key distinction: in token-level MDPs, each decision outputs one token; in turn-level MDPs, each decision produces a complete multi-token response before transitioning.
+At each turn $t$, the language model observes state $s_{t}$ and generates:
+$y_{t}=\langle\text{think}\rangle c_{t}\langle/\text{think}\rangle\langle\text{answer}\rangle a_{t}\langle/\text{answer}\rangle,$ (4)
+where $c_{t}$ externalizes reasoning and $a_{t}\in\mathcal{A}$ is the executable action. (See App.˜C for how existing SFT and RLVR paradigms adapt to turn-level MDPs.)
+C.2 Two-Player Zero-Sum Markov Games.
+We extend the single-agent MDP to competitive settings with a two-player zero-sum Markov game (Littman, 1994) $\mathcal{G}=(\mathcal{S},\mathcal{A}_{0},\mathcal{A}_{1},T,r,\gamma)$, where $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are the action spaces for player 0 and player 1 respectively.
+The zero-sum property requires:
+$r_{0}(s,a^{(0)},a^{(1)})+r_{1}(s,a^{(0)},a^{(1)})=0\quad\forall s,a^{(0)},a^{(1)},$ (5)
+where $a^{(0)}\in\mathcal{A}_{0}$ and $a^{(1)}\in\mathcal{A}_{1}$ denote actions taken by each player.
+Given trajectory $\tau=\{(s_{t},a_{t}^{(0)},a_{t}^{(1)})\}_{t=0}^{T}$, the returns satisfy $R_{1}(\tau)=-R_{0}(\tau)$.
+C.3 Supervised Fine-Tuning (SFT) in Turn-Level MDPs
 In the turn-level setting, SFT requires a dataset $\mathcal{D}_{\text{SFT}}=\{(s_{i},c_{i}^{*},a_{i}^{*})\}_{i=1}^{N}$ of states with expert reasoning traces $c_{i}^{*}$ and actions $a_{i}^{*}$.
@@
 The key limitation remains: SFT requires expensive human annotation of both reasoning traces and final answers.
-A.2 Reinforcement Learning with Verifiable Rewards (RLVR) in Turn-Level MDPs
+C.4 Reinforcement Learning with Verifiable Rewards (RLVR) in Turn-Level MDPs
 RLVR (DeepSeek Team, 2024) eliminates the need for reasoning supervision, requiring only state-answer pairs $\mathcal{D}_{\text{RLVR}}=\{(s_{i},a_{i}^{*})\}_{i=1}^{N}$. In the turn-level formulation:
@@
 where $r(s,a)=\mathbb{I}[a=a^{*}]$ indicates answer correctness and $y$ contains both reasoning and action.
-In single-turn settings without subsequent interactions, RLVR reduces to a contextual bandit problem (Langford and Zhang, 2007).
+In single-turn settings without subsequent interactions, RLVR reduces to a contextual bandit problem (Langford & Zhang, 2007).
 Recent works on mathematical (Shao et al., 2024; DeepSeek Team, 2024) and code reasoning (Zhu et al., 2024; Xin et al., 2024) show that even this simplified bandit-style RLVR can unlock sophisticated reasoning.
 However, these approaches still require human-curated problem sets $\mathcal{D}_{\text{RLVR}}$, which SPIRAL eliminates through self-play.
-Appendix B Experimental Setup Details
+Appendix D Experimental Setup Details
 This section provides complete implementation details for reproducing our experiments. We begin with visual examples of game environments, followed by our hyperparameter configurations.
-B.1 Game Environment Observations
+D.1 Game Environment Observations
 The language models receive structured text observations from each game environment.
-Fig.
-7 shows example observations from our three training games: TicTacToe, Kuhn Poker, and Simple Negotiation.
+Fig.˜7 shows example observations from our three training games: TicTacToe, Kuhn Poker, and Simple Negotiation.
 These observations serve as the input prompts $s_{t}$ at each turn, providing complete game state information in natural language format.
@@
 This ensures the model has sufficient information for decision-making despite hidden information.
-Similarly, Fig.
-8 presents observations from five evaluation environments used to test out-of-distribution generalization.
+Similarly, Fig.˜8 presents observations from five evaluation environments used to test out-of-distribution generalization.
 These games were never seen during training, allowing us to assess whether learned skills transfer to novel game mechanics.
 Figure 8: Example observations of five evaluation game environments.
-B.2 Hyperparameter Configuration
-Table 7 presents the complete hyperparameter settings used across all experiments.
+D.2 Hyperparameter Configuration
+Table˜6 presents the complete hyperparameter settings used across all experiments.
 These configurations were selected through preliminary experiments to balance training stability and computational efficiency.
@@
 Actor
-Maximum response length $4096$ tokens
+Maximum response length $8192$ tokens
 Sampling temperature 1.0
@@
 The distributed training infrastructure utilizes 8 H100 GPUs, with parallel actors generating game trajectories while a centralized learner performs synchronous policy updates.
-Appendix C Additional Results and Analysis
-C.1 Detailed Evolution of Case-by-Case Analysis
+On this hardware setup, the compute budget per experiment was approximately 25 hours for Qwen3-4B, 28 hours for Qwen3-8B, and 16 hours for both Llama3.1-8B-Instruct and Octothinker-8B.
+For SFT, we use Qwen3-32B (Yang et al., 2025) models to self-play on the targeted games, generating 25k winning trajectories and crafting them as the SFT dataset.
+D.3 Evaluation Settings
+To investigate whether the reasoning abilities developed through gameplay could transfer to non-game contexts, we evaluate our models on a suite of established benchmarks.
+All evaluations on these benchmarks are conducted in a zero-shot setting [^4] Except for the base model, for which we provide few-shot examples that follow the Qwen3 Report settings.
+to determine if game-induced reasoning could be successfully transferred to general problem-solving.
+We use a sampling temperature of $0.6$ and top-p of $0.95$ for all evaluations.
+Math Benchmarks.
+For mathematical reasoning, we use MATH500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME24, AIME25 (MAA, a), and AMC23 (MAA, b) datasets, which cover a wide range of topics including algebra, geometry, and competitive mathematics.
+Following the settings in Zhou et al.
+(2025), we report avg@32 for AIME24, AIME25 and AMC23; and pass@1 for other math benchmarks.
+General Reasoning Benchmarks.
+For general reasoning, we utilize GPQA-Diamond (Rein et al., 2024), which consists of graduate-level science questions, and MMLU-Pro (Wang et al., 2024), a benchmark for multidisciplinary knowledge.
+Appendix E Additional Results and Analysis
+E.1 Detailed Evolution of Case-by-Case Analysis
 To understand how reasoning patterns develop during training, we tracked the evolution of case-by-case analysis across checkpoints.
-Table 8 shows a concrete example from Minerva Math Problem 135, illustrating how models progressively develop structured reasoning.
-Table 8: Evolution of Case-by-Case Analysis in Minerva Problem 135
+Table˜7 shows a concrete example from Minerva Math Problem 135, illustrating how models progressively develop structured reasoning.
+Table 7: Case-by-Case Analysis in Minerva Problem 135
 Stage Problem: Round cosmic ages to nearest power of 10
@@
 Approach: Lumped reasoning without explicit case separation
-[…]
+<output> The age when helium formed is approximately 10 minutes… The age when hydrogen formed is 377,000 years… The current age is 13.8 Gyr.
+Rounding to nearest power of 10, we get: $\boxed{14}$ Gyr </output>
 Case Structure: No explicit cases - all subproblems handled in continuous text
@@
  1. Understanding the Context: BBN at $\sim$20 minutes
- 2. Nearest Power of 10: $1010^{1}=$ 10
-[…]
+ 2. Nearest Power of 10: $10^{1}=10$
+ 3. Final Answer: $\boxed{1}$ minute
 Case 2 - Subproblem 1:
  1. Understanding the Context: Recombination at $\sim$377,000 years
- 2. Nearest Power of 10: $00010^{5}=100,$ 000
-[…]
+ 2. Nearest Power of 10: $10^{5}=100,000$
+ 3. Final Answer: $\boxed{400000}$ years
 Case 3 - Subproblem 2:
@@
  2. Nearest Power of 10: Identifies $10^{10}$ correctly
-[…]
+ 3. Final Answer: $\boxed{13.8}$ Gyr </output>
 Case Structure: Clear case separation with consistent 3-step process
@@
 Approach: Systematic case-by-case with correct execution
-<think>
+<output>
 Case 1 - Subproblem 0:
@@
  2. Estimate Time: $\sim$20 minutes after Big Bang
-[…]
+ 3. Round to Power of 10: $\boxed{20}$ minutes
 Case 2 - Subproblem 1:
@@
  2. Estimate Time: $\sim$380,000 years
-[…]
+ 3. Round to Power of 10: $\boxed{400000}$ years
 Case 3 - Subproblem 2:
@@
  2. Convert Units: 13.8 billion years = 13.8 Gyr
-[…]
+ 3. Round to Power of 10: 13.8 $\rightarrow$ $^{1} = 10\boxed{10}$ Gyr </output>
 Case Structure: Complete systematic enumeration with correct logic
@@
 Early attempts show unorganized reasoning, while later checkpoints exhibit clear case separation and systematic analysis, a pattern that emerges from game playing and transfers to mathematical problem solving.
-C.2 Comprehensive Benchmark Results
-Table 9 presents extended results showing SPIRAL’s performance across different training configurations and base models.
-Table 9: SPIRAL training improves reasoning benchmarks for different base models
-Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average
+E.2 Extended Analysis of RAE Ablation Study
+Our ablation study of Role-conditioned Advantage Estimation reveals additional insights beyond those presented in Section 4 Figure 9 shows three additional metrics that further demonstrate RAE’s critical role in stable self-play training.
+Figure 9: Extended training dynamics comparing REINFORCE with RAE versus vanilla REINFORCE (continued from Figure 6).
+Left: Game win rates showing RAE achieves faster initial learning compared to vanilla REINFORCE.
+Middle: Math reasoning performance crashes from 35% to 12% without RAE.
+Right: Policy gradient norms exhibit instability then collapse to near-zero without RAE, while RAE maintains stable gradients around 0.1.
+Game Performance.
+The left panel shows that REINFORCE with RAE learns significantly faster, rapidly reaching 35% win rate while vanilla REINFORCE exhibits slower initial learning.
+RAE’s superior learning efficiency demonstrates the benefits of role-conditioned advantage estimation for accelerating policy improvement.
+Math Reasoning Collapse.
+The middle panel reveals the most dramatic failure: math reasoning performance without RAE crashes from 35% to 12% at around step 150 (a 66% relative decrease).
+This collapse occurs precisely when models stop generating reasoning traces, confirming that thinking collapse directly causes reasoning failure.
+Gradient Stability.
+The right panel shows policy gradient norms, revealing the underlying optimization dynamics.
+Without RAE, gradients exhibit high variance with erratic spikes before collapsing to near-zero after step 200, indicating convergence to a degenerate policy.
+RAE maintains stable gradient norms around 0.1 throughout training, enabling continuous improvement.
+These additional results reinforce our main finding: self-play alone is insufficient for stable reasoning development.
+Proper variance reduction through role-specific baselines is essential to prevent models from converging to degenerate policies that abandon reasoning in favor of minimal outputs.
+E.3 Game Generalization
+We conducted additional experiments on higher-complexity variants of our training games (5x5 Tic-Tac-Toe, 5-Card Kuhn Poker, and 8-Resource Negotiation) to assess scalability.
+Table 8: Generalization performance on increased complexity environments. SPIRAL retains significantly higher performance on out-of-distribution (OOD) tasks compared to SFT.
+Environment Setting Qwen3-4B Base SFT SPIRAL (Ours)
+Tic-Tac-Toe $\times3$ (Train) 17.5 46.9 54.3
+$\times5$ (OOD) 5.5 17.8 27.3
+Kuhn Poker 3-Card (Train) 21.5 43.6 53.9
+5-Card (OOD) 21.9 28.6 50.1
+Simple Negotiation 5 Resources (Train) 15.6 26.7 33.2
+8 Resources (OOD) 5.5 8.7 31.0
+Average (Train) 18.2 39.1 47.1
+Average (OOD) 11.0 18.4 36.1
+Table 8 shows that although SFT has comparable improvement over the base model in in-domain environments (averaging $39.1\%$ vs.
+$18.2\%$), it struggles to adapt to OOD cases with increased complexity.
+The SFT model’s performance drops to only $18.4\%$.
+In contrast, SPIRAL demonstrates better generalizability.
+It not only outperforms SFT in the training settings (averaging $47.1\%$) but maintains robust performance in the more complex OOD environments, achieving an average of $36.1\%$.
+This suggests that SPIRAL is more capable of handling increased complexity.
+E.4 Comprehensive Benchmark Results
+Table˜9 presents extended results showing SPIRAL’s performance across different training configurations and base models.
+Table 9: SPIRAL training improves reasoning benchmarks for different base models.
+We include an additional baseline of SFT on 52k trajectories to demonstrate that simply scaling supervised data does not match the gains from self-play.
+Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA-D MMLU-Pro Average
 Qwen3-4B-Base Family
-Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 33.1
-+ SFT 76.0 16.7 13.3 38.4 50.0 31.2 33.0 48.8 38.4
-+ Mistral Opponent (KuhnPoker) 64.0 6.7 6.7 29.8 50.0 26.1 35.6 43.6 32.8
-+ Gemini Opponent (KuhnPoker) 69.2 10.0 13.3 33.8 50.0 33.8 35.3 55.5 37.6
+Qwen3-4B-Base 73.4 9.6 6.2 33.3 42.4 29.4 30.6 47.2 34.0
++ SFT (Kuhn) 74.0 11.0 10.4 36.7 48.6 36.8 33.0 48.8 37.4
++ SFT (Multi, 25k) 74.2 13.7 11.7 37.6 51.1 40.1 37.8 51.3 39.7
++ SFT (Multi, 52k) 73.4 16.2 11.9 39.9 48.7 38.7 38.2 51.2 39.7
++ Mistral Opponent (KuhnPoker) 64.0 4.3 2.1 29.8 31.6 26.1 35.6 43.6 29.6
++ Gemini Opponent (KuhnPoker) 69.2 5.2 4.7 33.8 29.8 33.8 35.3 55.5 33.4
 + SPIRAL (TicTacToe) 75.6 10.0 13.3 38.5 55.0 42.6 37.6 57.7 41.3
-+ SPIRAL (KuhnPoker) 76.4 13.3 10.0 38.4 57.5 42.4 37.0 56.5 41.4
-+ SPIRAL (Negotiation) 76.4 10.0 10.0 41.2 57.5 36.0 34.7 54.2 40.0
-+ SPIRAL (TicTacToe+KuhnPoker) 76.2 13.3 16.7 40.7 60.0 41.5 35.7 57.2 42.7
-+ SPIRAL (TicTacToe+KuhnPoker+Negotiation) 74.2 20.0 16.7 39.4 55.0 38.6 36.9 57.7 42.3
++ SPIRAL (KuhnPoker) 76.4 18.2 15.6 38.4 61.2 42.4 37.0 57.7 43.4
++ SPIRAL (Negotiation) 75.6 11.7 10.2 38.1 51.7 39.3 36.7 57.0 40.0
++ SPIRAL (TicTacToe+KuhnPoker) 76.2 11.4 10.7 40.7 57.2 41.5 35.7 57.2 41.3
++ SPIRAL (Multi-Game) 78.2 19.7 13.3 41.8 61.6 42.6 40.1 58.5 44.5
+Qwen3-8B-Base Family
+Qwen3-8B-Base 77.0 12.1 11.2 33.5 50.6 38.2 38.0 55.7 39.5
++ SFT (Multi, 25k) 82.8 19.9 15.6 45.9 63.5 40.8 41.6 58.8 46.1
++ SFT (Multi, 52k) 81.8 22.6 14.0 48.8 61.6 40.0 43.1 57.4 46.2
++ SPIRAL (Multi-Game) 86.6 26.2 16.8 49.6 65.2 46.3 44.6 61.1 49.6
+Octothinker-8B-Base Family
+Octothinker-8B-Base 65.6 1.7 0.5 26.6 33.5 25.7 22.1 30.8 25.8
++ SFT (Multi, 25k) 66.0 3.3 3.8 23.9 31.0 23.8 24.9 39.1 27.0
++ SFT (Multi, 52k) 66.4 4.2 3.6 21.9 31.6 24.6 24.9 40.8 27.3
++ SPIRAL (Multi-Game) 68.6 5.3 4.8 33.7 43.2 32.0 33.8 49.3 33.8
+Llama-3.1-8B-Instruct Family
+Llama-3.1-8B-Instruct 46.4 4.6 0.7 13.8 23.3 22.8 30.2 49.1 23.9
++ SFT (Multi, 25k) 51.8 4.6 0.7 19.1 23.3 21.7 30.0 48.9 25.0
++ SFT (Multi, 52k) 51.0 4.3 0.0 21.5 23.2 21.1 30.2 49.1 25.1
++ SPIRAL (Multi-Game) 49.8 4.9 1.8 17.3 26.0 24.6 32.2 50.4 25.9
 DeepSeek-Distill-Qwen-7B Family
-DeepSeek-Distill-Qwen-7B 90.8 46.7 36.7 56.9 92.5 48.2 48.6 57.1 59.7
-+ SPIRAL (Multi-Game) 93.0 43.3 46.7 57.9 92.5 51.1 49.6 58.9 61.7
+DeepSeek-Distill-Qwen-7B 90.8 53.0 39.5 56.9 89.3 48.2 48.6 57.1 60.4
++ SFT (Multi) 91.8 49.3 36.6 52.4 88.2 48.2 44.5 55.6 58.3
++ SPIRAL (Multi-Game) 93.0 54.1 40.8 57.9 89.3 51.1 49.6 58.9 61.8
 These results reveal several important insights.
@@
 Third, SPIRAL improves even strong models like DeepSeek-Distill-Qwen-7B (from 59.7% to 61.7%), demonstrating that competitive game self-play training can enhance models that already excel at reasoning tasks.
-Appendix D Case Study Methodology
+We curated an additional 27k dataset with Qwen3-32B self-play and evaluated SFT on a total of 52k trajectories but training with 1 epoch.
+As shown in the table, doubling the SFT data yields no major improvement (e.g., Qwen3-4B Average 39.7% vs 39.7%), while SPIRAL consistently outperforms both SFT baselines.
+This confirms that the benefits of SPIRAL stem from the reinforcement learning dynamic rather than simply dataset size.
+E.5 Statistical Robustness
+To address concerns regarding statistical significance, we re-ran our main experiments with 3 random seeds (seeds 14, 42, 100).
+As shown in Table˜10, SPIRAL consistently outperforms the SFT baseline with narrow confidence intervals, confirming the robustness of our gains.
+Table 10: Comprehensive performance comparison across multiple seeds. SPIRAL demonstrates consistent improvement over baselines across diverse benchmarks.
+Model Math500 AIME24 AIME25 Olympiad AMC-23 Minerva GPQA-D MMLU-Pro Average
+Qwen3-4B-Base 73.4 9.6 6.2 33.3 42.4 29.4 30.6 47.2 34.0
++ SFT-Multi $74.0\pm 0.6$ $2.4\pm 1.5 11$$1.2\pm 1.1$ $37.8\pm 0.3$ $52.2\pm 0.8$ $40.7\pm 0.5$ $37.7\pm 1.2$ $50.9\pm 0.5$ $39.6\pm 0.4$
++ SPIRAL-Multi (Ours) $\mathbf{78.7\pm 2.0}$ $\mathbf{18.8\pm 2.5}$ $\mathbf{15.0\pm 1.3}$ $\mathbf{41.8\pm 1.3}$ $\mathbf{62.0\pm 1.6}$ $\mathbf{42.1\pm 1.3}$ $\mathbf{39.1\pm 3.1}$ $\mathbf{58.4\pm 0.5}$ $\mathbf{44.5\pm 0.5}$
+E.6 Game Trajectory Statistics
+In Table˜11, we added a comprehensive table with the average number of game lengths, reasoning tokens per step, P1/P2 self-play win rates, and average win-rate and win-rate per game against Gemini-2.0-Flash.
+Table 11: Game trajectory statistics across training checkpoints. We observe increasing game length and reasoning tokens alongside improved win rates against the fixed opponent.
+Training Avg. Game Round Avg. Reasoning Self-Play Self-Play Avg. Win-Rate Win-Rate vs. Gemini Win-Rate vs. Gemini Win-Rate vs. Gemini
+Steps (Moves) Tokens / Step P1 Win-Rate P2 Win-Rate vs. Gemini (TicTacToe) (Kuhn Poker) (Simple Negotiation)
+Step 0 1.69 4061 42.7% 57.3% 12.5% 12.5% 6.25% 18.8%
+Step 128 7.63 1609 48.7% 51.3% 28.5% 16.7% 37.5% 31.2%
+Step 256 8.47 1755 53.2% 46.8% 51.8% 42.9% 68.8% 43.8%
+Step 384 9.43 1921 59.1% 40.9% 66.7% 75.0% 68.8% 56.3%
+Step 400 9.55 2032 62.4% 37.6% 67.4% 83.3% 62.5% 56.3%
+E.7 SPIRAL as part of the mid-training stage
+[…]
+Table 12: Performance comparison of SPIRAL integrated into different training stages. SPIRAL acts as a robust performance booster both before and after RLVR.
+Model Math500 AIME24 AIME25 Olympiad AMC-23 Minerva GPQA-D MMLU-Pro Average
+Qwen3-4B-Base 73.4 9.6 6.2 33.3 42.4 29.4 30.6 47.2 34.0
+RLVR (Math) 83.0 18.4 15.6 44.6 62.8 43.4 43.7 56.8 46.0
+SPIRAL 78.2 19.7 13.3 41.8 61.6 42.6 40.1 58.5 44.5
+SPIRAL $\to$ RLVR 84.2 23.1 17.2 45.2 59.6 42.1 43.2 57.4 46.5
+RLVR $\to$ SPIRAL 86.1 22.6 18.1 46.0 62.5 44.3 44.7 60.8 47.9
+Our results in Table˜12 validate SPIRAL’s effectiveness on integrating as part of the mid-training: SPIRAL $\to$ RLVR outperforms the Base $\to$ RLVR baseline, while applying SPIRAL as a post-RLVR stage (RLVR $\to$ SPIRAL) yields the highest performance (Avg 48.1).
+Furthermore, on the stronger DeepSeek-Distill-Qwen-7B, SPIRAL-Multi improves the average score to 59.3, whereas standard SFT leads to performance regression.
+These findings demonstrate that SPIRAL serves as a robust booster that mitigates the alignment tax often seen in standard fine-tuning stages.
+E.8 Instruct Model Results
+To better study our method’s effectiveness, we perform additional experiments with Qwen3-4B-Instruct-2507.
+As shown in the Table˜13, standard SFT on reasoning trajectories slightly degrades performance.
+In contrast, SPIRAL successfully reverses this trend.
+SPIRAL achieves a 2% improvement over the original model and a substantial average gain of 3.6% over the SFT baseline, reaching a total average score of 75.51%.
+This demonstrates that self-play can refine genuine reasoning capabilities even in models that have strong performance.
+Table 13: Performance comparison on Qwen3-4B-Instruct. While standard SFT leads to regression, SPIRAL improves reasoning capabilities across most benchmarks.
+Model Math500 AIME24 AIME25 Olympiad AMC-23 Minerva GPQA-D MMLU-Pro Average
+Qwen3-4B-Instruct (Base) 91.2 64.6 47.4 89.6 82.1 87.6 62.0 69.6 74.10
++ SFT-Multi 90.8 59.3 46.1 86.4 79.6 82.3 61.4 69.1 71.88
++ SPIRAL-Multi (Ours) 93.5 67.9 49.8 88.8 83.3 85.5 61.8 71.5 75.91
+Appendix F Case Study Methodology
 This section details our systematic approach to discovering and analyzing reasoning pattern transfer from games to mathematics.
 Rather than searching for predetermined patterns, we employed a bottom-up discovery process to identify what reasoning strategies naturally emerge and transfer between domains.
-D.1 Data Collection Framework
+F.1 Data Collection Framework
 Our analysis examined reasoning traces from two sources across three training checkpoints:
@@
 Temporal Analysis: Checkpoints at steps 0 (initial), 128 (intermediate), and 400 (final) capture the evolution of reasoning complexity throughout training.
-D.2 Bottom-Up Pattern Discovery Process
+F.2 Bottom-Up Pattern Discovery Process
 Rather than searching for predefined patterns, we employed GPT-4.1 to discover patterns that naturally emerge in the data.
@@
     Pattern Recognition: Identifying regularities and structures
-D.3 Cross-Domain Transfer Quantification
+F.3 Cross-Domain Transfer Quantification
 After discovering patterns in each domain, we compared them to identify which strategies transfer between games and mathematics:
@@
 Mathematics inherently requires pattern identification, so game training enhances an already-essential mathematical skill, producing stronger pattern recognition than games alone develop.
-D.4 Pattern Evolution Analysis
+F.4 Pattern Evolution Analysis
 To understand how reasoning develops during training, we tracked pattern emergence across checkpoints:
@@
 Show concrete examples of improvement Focus on the actual evolution you can see in the data.
-D.5 Concrete Transfer Example Identification
+F.5 Concrete Transfer Example Identification
 To validate transfer claims, we identified parallel reasoning structures across domains:
@@
 Rate the clarity of transfer (1-10) Focus on examples where the same reasoning approach clearly appears in both domains.
-D.6 Pattern Classification at Scale
+F.6 Pattern Classification at Scale
 After discovering patterns through bottom-up analysis, we classified all traces to measure transfer rates:
@@
 Be strict - only include clear examples of each pattern.
-D.7 Validation Methodology
+F.7 Validation Methodology
 To ensure robust findings, we implemented multiple validation steps:
@@
 This methodology ensures our findings reflect genuine cognitive transfer rather than superficial pattern matching, providing quantitative evidence that competitive gameplay develops reasoning skills applicable far beyond the training domain.
-Appendix E Game Environment Specifications
+Appendix G Game Environment Specifications
 This section provides detailed specifications for all game environments used in our experiments, including both training and evaluation games.
-E.1 Training Game Environments
+G.1 Training Game Environments
 TicTacToe tests spatial pattern recognition through perfect information gameplay.
@@
 These capabilities should transfer to probability problems and strategic decision-making.
-Simple Negotiation develops multi-constraint optimization through resource trading.
-Two players exchange Wood and Gold tokens, with opposing utility functions creating natural tension.
-Each player aims to maximize portfolio value through proposals and counteroffers.
-Success requires understanding opponent preferences, planning multi-step trades, and strategic communication.
-We expect these skills to enhance optimization problems and multi-constraint reasoning tasks.
-E.2 Out-of-Distribution Evaluation Games
+Simple Negotiation is a game that develops multi-constraint optimization skills through resource trading.
+Two players exchange Wheat, Wood, Sheep, Brick, and Gold tokens.
+Since the utility of these resources varies for each player, there is a natural incentive to trade.
+Each player aims to maximize their portfolio’s value by making proposals and counteroffers.
+Success requires understanding an opponent’s preferences, planning multi-step trades, and communicating strategically.
+We expect these skills to improve performance on optimization problems and multi-constraint reasoning tasks.
+G.2 Out-of-Distribution Evaluation Games
 Our evaluation suite tests whether learned skills generalize to novel mechanics:
@@
 This tests whether static pattern recognition from TicTacToe transfers to trajectory planning and dynamic obstacle avoidance.
-Connect Four adds directional constraints to spatial reasoning.
-Players drop discs into columns with gravity, aiming to connect four vertically, horizontally, or diagonally.
-The vertical placement mechanic tests whether TicTacToe strategies generalize to games with restricted movement options.
 Pig Dice isolates risk-reward decision making.
@@
 This tests whether probabilistic reasoning from Kuhn Poker extends to sequential risk assessment and expected value calculation in different contexts.
-Liar’s Dice combines probability with deception detection.
-Players make escalating claims about hidden dice totals and challenge suspected bluffs.
-Success requires probability calculation, opponent modeling, and strategic deception, testing whether Kuhn Poker skills scale to multi-round bluffing games.
 Truth and Deception focuses on asymmetric information and persuasion.